### v4

# Residually Constructible Extensions

Abstract.
Let $T$ be an o-minimal theory expanding $\mathrm{RCF}$ and $T_{\mathrm{convex}}$ be the common theory of its models expanded by predicate for a non-trivial $T$-convex valuation ring. We call an elementary extension $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{*},\mathcal{O}_{*})\models T_{\mathrm{convex}}$ res-constructible if there is a tuple $\overline{s}$ in $\mathcal{O}_{*}$ such that the projection $\operatorname{\mathbf{res}}(\overline{s})$ of $\overline{s}$ in the residue field sort is $\mathrm{dcl}$-independent over the residue field $\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$ of $(\mathbb{E},\mathcal{O})$ and such that $\operatorname{\mathrm{dcl}}(\mathbb{E},\overline{s})=\mathbb{E}_{*}$. We study factorization properties of res-constructible extensions. Our main result is that a res-constructible extension $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{*},\mathcal{O}_{*})$ has the property that all $(\mathbb{E}_{1},\mathcal{O}_{1})$ with $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{1},\mathcal{O}_{1})\prec(\mathbb{E}_{*},\mathcal{O}_{*})$ are res-constructible over $(\mathbb{E},\mathcal{O})$, if and only if $\mathbb{E}_{*}$ has countable dimension over $\mathbb{E}$ or the value group $\operatorname{\mathbf{v}}(\mathbb{E}_{*},\mathcal{O}_{*})$ contains no uncountable well-ordered subset. This analysis entails complete answers to [9, Problem 5.12].
Key words and phrases:
model theory, o-minimality, valued fields, residually constructible extensions, pseudocompletions, weakly immediate types
2020 Mathematics Subject Classification:
Primary 03C64; Secondary 12J10 12J15
The second author’s work was supported by EPSRC DTP 2224 [EP/W524372/1]
1. Introduction
Let $T$ be a complete o-minimal $L$-theory expanding the theory $\mathrm{RCF}$ of real closed fields in some language $L$ containing the language of ordered rings. There has been significant work on this class of theories; for example $T$ could be the theory $T_{\exp}$ of the field of reals expanded by the natural exponential function ([11]) or the theory $T_{an,\exp}$ of the reals expanded by the natural exponential and all restricted analytic functions ([4]).
Recall from [3, (2.7)], that a $T$-convex valuation subring $\mathcal{O}$ of a model $\mathbb{E}\models T$ is a convex subring closed under all continuous $\emptyset$-definable functions $f:\mathbb{E}\to\mathbb{E}$.
By results of van den Dries and Lewenberg [3, (3.10), (3.13), (3.14)], the common theory $T_{\mathrm{convex}}$ in the language $L_{\mathrm{convex}}:=L\cup\{\mathcal{O}\}$ of all models of $T$ expanded by a unary predicate $\mathcal{O}$ interpreted as a non-trivial $T$-convex valuation ring, is complete and weakly o-minimal.
Results of [3], [1], and [10] entail that if $T$ is power bounded with field of exponents $\Lambda$, then for every elementary extension $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})\models T_{\mathrm{convex}}$ one can find a $\operatorname{\mathrm{dcl}}_{T}$-basis of $\mathbb{E}_{*}$ over $\mathbb{E}$ of the form $(\overline{r},\overline{v},\overline{e})$ where:
  1. (1)
    $\overline{r}$ is a tuple in $\mathcal{O}_{*}$ such that the projection $\operatorname{\mathbf{res}}(\overline{r})$ of $\overline{r}$ in the residue field $\operatorname{\mathbf{res}}(\mathbb{E}_{*})$ is $\operatorname{\mathrm{dcl}}_{T}$-independent over the residue field $\operatorname{\mathbf{res}}(\mathbb{E})$ of $(\mathbb{E},\mathcal{O})$;
  2. (2)
    $\overline{v}$ is a tuple in $\mathbb{E}_{*}$ such that $\operatorname{\mathbf{v}}(\overline{v})$ is $\Lambda$-linearly independent over the value group $\operatorname{\mathbf{v}}(\mathbb{E})$ of $\mathbb{E}$;
  3. (3)
    $(\mathbb{E}_{*},\mathcal{O}_{*})$ is an immediate extension of $(\mathbb{E}\langle\overline{r},\overline{v}\rangle,\mathcal{O}_{*}\cap\mathbb{E}\langle\overline{r},\overline{v}\rangle)$, where $\mathbb{E}\langle\overline{r},\overline{v}\rangle:=\operatorname{\mathrm{dcl}}(\mathbb{E},\overline{r},\overline{v})$.
Even when $T$ is not power bounded, we will call tuples as $\overline{r}$ in item (1) above $\mathcal{O}_{*}$-res-constructions over $\mathbb{E}$ (Definition 14). In this paper we will be concerned with res-constructible elementary extensions of models of $T_{\mathrm{convex}}$, that is, elementary extensions $(\mathbb{E}_{*},\mathcal{O}_{*})\succeq(\mathbb{E},\mathcal{O})$ such that for some $\mathcal{O}_{*}$-res-construction $\overline{r}$ over $\mathbb{E}$, $\mathbb{E}_{*}=\mathbb{E}\langle\overline{r}\rangle$.
Specific instances of res-constructible extensions were considered in [9]: more precisely a res-constructible extension $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})$ such that the underlying residue field extension $\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})\preceq\operatorname{\mathbf{res}}(\mathbb{E}_{*},\mathcal{O}_{*})$ is a Cauchy-completion, is exactly a pseudo-completion of $\mathbb{E}$ with respect to $\{\mathcal{O}\}$ in the sense of [9, Def. 4.2].
After showing that pseudo-completions are unique up to isomorphisms but admit non-surjective self-embeddings, Tressl in [9, Problem 5.12] asks whether given $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}^{\prime},\mathcal{O}^{\prime})\preceq(\mathbb{E}_{1},\mathcal{O}_{1})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})$ such that $\mathbb{E}_{*}$ and $\mathbb{E}^{\prime}$ are pseudo-completions of $\mathbb{E}$ with respect to $\{\mathcal{O}\}$, one has that necessarily $\mathbb{E}_{1}$ is also a pseudo-completion of $\mathbb{E}$ with respect to $\{\mathcal{O}\}$.
In this paper we will study the more general question of whether res-constructible extensions are closed under taking right-factors, i.e. whether (or rather more appropriately “when”) given a composition of elementary extensions $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{1},\mathcal{O}_{1})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})$ of models of $T$-convex such that the composite extension $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})$ is res-constructible, one can deduce that $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{1},\mathcal{O}_{1})$ is also res-constructible.
We show that in general res-constructible extensions are are not closed under right factors and give a complete characterization of those res-constructible extensions $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})$ all of whose right-factors $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{1},\mathcal{O}_{1})$ are res-constructible.
Theorem A.
Suppose $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})$ is res-constructible. Then the following are equivalent:
  1. (1)
    $\dim_{\operatorname{\mathrm{dcl}}}(\mathbb{E}_{*}/\mathbb{E})$ is countable or the value group $\operatorname{\mathbf{v}}(\mathbb{E}_{*},\mathcal{O}_{*})$ of $(\mathbb{E}_{*},\mathcal{O}_{*})$ contains no uncountable ordinal;
  2. (2)
    for all $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{1},\mathcal{O}_{1})\prec(\mathbb{E}_{*},\mathcal{O}_{*})$, the extension $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{1},\mathcal{O}_{1})$ is res-constructible;
  3. (3)
    for all $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{1},\mathcal{O}_{1})\prec(\mathbb{E}_{*},\mathcal{O}_{*})$, with $\operatorname{\mathbf{res}}(\mathbb{E}_{1},\mathcal{O}_{1})=\operatorname{\mathbf{res}}(\mathbb{E}_{*},\mathcal{O}_{*})$, the extension $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{1},\mathcal{O}_{1})$ is res-constructible.
This implies that the answer to [9, Problem 5.12] is affirmative if and only if the value group $\operatorname{\mathbf{v}}(\mathbb{E}_{*},\mathcal{O}_{*})$ does not contain uncountable ordinals or $\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$ has a Cauchy-completion of countable $\operatorname{\mathrm{dcl}}$-dimension (Section 4).
We will now spend some words about the proof of the main implication $(1)\Rightarrow(2)$ of Theorem A.
It is not to hard to show that if $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})$ is res-constructible, then for any intermediate $\mathbb{E}\preceq\mathbb{E}_{1}\preceq\mathbb{E}_{*}$, such that $\dim_{\operatorname{\mathrm{dcl}}}(\mathbb{E}_{1}/\mathbb{E})<\aleph_{0}$, one has that $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{1},\mathcal{O}_{*}\cap\mathbb{E}_{1})$ and $(\mathbb{E}_{1},\mathcal{O}_{*}\cap\mathbb{E}_{1})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})$ are both res-constructible extensions (Corollary 21). This easily entails that even if $\dim_{\operatorname{\mathrm{dcl}}}(\mathbb{E}_{1}/\mathbb{E})=\aleph_{0}$, one has that $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{1},\mathcal{O}_{*}\cap\mathbb{E}_{1})$ is res-constructible however it is also not hard to show that $(\mathbb{E}_{1},\mathcal{O}_{*}\cap\mathbb{E}_{1})\prec(\mathbb{E}_{*},\mathcal{O}_{*})$ doesn’t need to be res-constructible anymore.
We will prove $(1)\Rightarrow(2)$ of Theorem A by building a $\mathcal{O}_{*}$-res-construction $\overline{s}$ for $\mathbb{E}_{1}$ over $\mathbb{E}$ as an ordinal indexed sequence $\overline{s}:=(s_{i}:i<\lambda)$. By the already mentioned Corollary 21 about factors with finite dimension, the main problem in the transfinite construction of $\overline{s}$ is at limit steps: in fact it may happen that even if $\mathbb{E}_{1}$ was res-constructible over $\mathbb{E}$, for some $\mathcal{O}_{*}$-res-construction, say $(s_{i}:i<\omega)$, it is not anymore res-constructible over over $\mathbb{E}\langle s_{i}:i<\omega\rangle$. This difficulty will be overcome by setting up the induction so that at each limit step the res-construction $(s_{i}:i<\omega\cdot\alpha)$ satisfies both a model theoretic orthogonality condition and a closure condition (Definition 25) relative to a given $\mathcal{O}_{*}$-res-construction $\overline{r}$ of $\mathbb{E}_{*}$ over $\mathbb{E}$.
Subsequent to proving Theorem A, in Section  4, we will turn to [9, Problem 5.12], and give a complete answer, dependent on the properties of $\mathbb{E}_{*}$. We will also show that in the power-bounded setting, this dependence can be reduced to considering properties of $\mathbb{E}$.
Acknowledgments
[…]
2. Preliminaries and notation
In the following $T\supseteq\mathrm{RCF}$ will be some fixed complete o-minimal theory in some language $L$ containing the language of ordered rings. If $\mathbb{E}\models T$ and $x$ is a tuple or a set in some $\mathbb{U}\succ\mathbb{E}$, we will denote the definable closure $\operatorname{\mathrm{dcl}}_{T}(\mathbb{E}\cup x)$ by $\mathbb{E}\langle x\rangle$. Recall that this is still a model of $T$ because $T$ has definable Skolem functions. Also, for a subset $S\subseteq\mathbb{E}$, we will denote the convex hull of $S$ in $\mathbb{U}$ by $\operatorname{\mathrm{CH}}_{\mathbb{U}}(S):=\{x\in\mathbb{U}:\exists s_{1},s_{2}\in S,\;s_{1}\leq x\leq s_{2}\}$. Finally, if $R\subseteq X\times X$ is a binary relation on a set $X$, given subsets $A,B\subseteq X$, we will write $A^{RB}$ for the set $A^{RB}=\{x\in A:\forall b\in B,\,xRb\}$; if $B=\{b\}$ is a singleton we will write $A^{Rb}$ for $A^{R\{b\}}$. Usually $R$ will be an order relation or equality.
2.1. Valuation-related notation.
Given a field $\mathbb{E}$ and a valuation ring $\mathcal{O}$ on $\mathbb{E}$, we will denote by $\mathfrak{o}$ the unique maximal ideal of $\mathbb{E}$. The value group will be denoted by $(\operatorname{\mathbf{v}}(\mathbb{E},\mathcal{O}),+):=(\mathbb{E}^{\neq 0},\cdot)/(\mathcal{O}\setminus\mathfrak{o})$ or by $\operatorname{\mathbf{v}}_{\mathcal{O}}\mathbb{E}$. The valuation will be the quotient $\operatorname{\mathbf{v}}_{\mathcal{O}}:\mathbb{E}^{\neq 0}\to\operatorname{\mathbf{v}}(\mathbb{E},\mathcal{O})$, and will be extended to the whole $\mathbb{E}$ by setting $\operatorname{\mathbf{v}}(0)=\infty$. The value group will be ordered as usual by setting $\operatorname{\mathbf{v}}(\mathbb{E},\mathcal{O})^{\geq 0}:=\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathcal{O}^{\neq 0})$.
The associated dominance, weak asymptotic equivalence, strict dominance, and strict asymptotic equivalence on $\mathbb{E}$ will be denoted respectively by:
$x\preceq y$ $\Longleftrightarrow\operatorname{\mathbf{v}}_{\mathcal{O}}(x)\geq\operatorname{\mathbf{v}}_{\mathcal{O}}(y)\Leftrightarrow x/y\in\mathcal{O};$
$x\asymp y$ $\Longleftrightarrow\operatorname{\mathbf{v}}_{\mathcal{O}}(x)=\operatorname{\mathbf{v}}_{\mathcal{O}}(y)\Leftrightarrow x/y\in\mathcal{O}\setminus\mathfrak{o};$
$x\prec y$ $\Longleftrightarrow\operatorname{\mathbf{v}}_{\mathcal{O}}(x)>\operatorname{\mathbf{v}}_{\mathcal{O}}(y)\Leftrightarrow x/y\in\mathfrak{o};$
$x\sim y$ $\Longleftrightarrow x-y\prec x\Leftrightarrow 1-x/y\in\mathfrak{o}.$
The imaginary sort given by the quotient under the strict asymptotic equivalence $\sim$ will be called rv-sort and denoted by $\operatorname{\mathbf{rv}}(\mathbb{E},\mathcal{O}):=(\mathbb{E}^{\neq 0},\cdot)/(1+\mathfrak{o})$ or $\operatorname{\mathbf{rv}}_{\mathcal{O}}(\mathbb{E})$, the quotient map will be denoted by $\operatorname{\mathbf{rv}}_{\mathcal{O}}:\mathbb{E}^{\neq 0}\to\operatorname{\mathbf{rv}}(\mathbb{E},\mathcal{O})$.
Finally the residue field sort will be denoted by $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})=\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O}):=\mathcal{O}/\mathfrak{o}$ and the corresponding quotient map by $\operatorname{\mathbf{res}}_{\mathcal{O}}:\mathcal{O}\to\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$. We extend $\operatorname{\mathbf{res}}_{\mathcal{O}}$ to a total function $\operatorname{\mathbf{res}}_{\mathcal{O}}:\mathbb{E}\to\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$ by setting $\operatorname{\mathbf{res}}_{\mathcal{O}}$ to be 0 outside of $\mathcal{O}$.
The subscript $\mathcal{O}$ will be omitted when this does not create ambiguity.
2.2. Main facts on $T$-convex valuations
Recall that if $\mathbb{E}\models T$, a subset $\mathcal{O}\subseteq\mathbb{E}$ is said to be a $T$-convex subring when it is order-convex and closed under $\emptyset$-definable continuous total unary functions ([3]).
Fact 1 (van den Dries and Lewenberg [3]).
Let $\mathbb{U}\succeq\mathbb{E}\models T$ and let $\mathcal{O}^{\prime}\subseteq\mathbb{U}$ and $\mathcal{O}\subseteq\mathbb{E}$ be $T$-convex subrings with $\mathcal{O}^{\prime}\cap\mathbb{E}=\mathcal{O}$. If $x\in\mathbb{U}\setminus\mathbb{E}$, then $\mathcal{O}^{\prime}\cap\mathbb{E}\langle x\rangle\in\{\mathcal{O}_{x},\mathcal{O}_{x}^{-}\}$ where
$\mathcal{O}_{x}:=\{y\in\mathbb{E}\langle x\rangle:y<\mathbb{E}^{>\mathcal{O}}\}\quad\text{and}\quad\mathcal{O}_{x}^{-}:=\operatorname{\mathrm{CH}}_{\mathbb{E}\langle x\rangle}(\mathcal{O}).$
Definition 2.
Let $L_{\mathrm{convex}}:=L\cup\{\mathcal{O}\}$ be the language $L$ expanded by a unary predicate $\mathcal{O}$ and let $T_{\mathrm{convex}}^{-}$ be the common theory in $L_{\mathrm{convex}}$ of all expansions $(\mathbb{E},\mathcal{O})$ of models of $T$ where $\mathcal{O}$ is interpreted as a $T$-convex subring. Let $T_{\mathrm{convex}}$ be the theory $T_{\mathrm{convex}}^{-}\cup\{\exists x\notin\mathcal{O}\}$.
Fact 3 (van den Dries and Lewenberg [3]).
The theory $T_{\mathrm{convex}}$ is complete and weakly o-minimal. Moreover if $T$ is universally axiomatized and eliminates quantifiers, then $T_{\mathrm{convex}}$ eliminates quantifiers.
Remark 4.
Notice that if $T^{\mathrm{def}}$ is the expansion by definition of $T$ to the language $L^{\mathrm{def}}$ with a symbol for every $T$-definable function, then $T$ is model complete (in fact it is universally axiomatized and eliminates quantifiers). In particular Fact 3 implies that if $(\mathbb{E},\mathcal{O})\models T_{\mathrm{convex}}$ and $\mathbb{E}\succeq\mathbb{E}^{\prime}\models T$ is such that $\mathbb{E}^{\prime}\not\subseteq\mathcal{O}$, then $(\mathbb{E},\mathcal{O})\succeq(\mathbb{E}^{\prime},\mathcal{O}\cap\mathbb{E}^{\prime})$.
Definition 5.
An elementary substructure $\mathbb{K}\preceq\mathbb{E}\models T$ is said to be tame or definably Dedekind-complete (denoted $\mathbb{K}\preceq_{\mathrm{tame}}\mathbb{E}$) if for all $\mathbb{E}$-definable $S\subseteq\mathbb{E}$, if $S\cap\mathbb{K}$ is bounded in $\mathbb{K}$, then it has a supremum in $\mathbb{K}$. If $\mathbb{K}\preceq_{\mathrm{tame}}\mathbb{E}$, then there is a unique function $\operatorname{\mathrm{st}}_{\mathbb{K}}:\operatorname{\mathrm{CH}}_{\mathbb{E}}(\mathbb{K})\to\mathbb{K}$ with the property that $\operatorname{\mathrm{st}}_{\mathbb{K}}(x)-x<\mathbb{K}^{>0}$ for all $x\in\operatorname{\mathrm{CH}}_{\mathbb{E}}(\mathbb{K})$.
Fact 6 (Marker and Steinhorn [7]).
$\mathbb{K}\preceq_{\mathrm{tame}}\mathbb{E}$ if and only if whenever $c$ is a tuple of $\mathbb{E}$ then $\operatorname{\mathrm{tp}}(c/\mathbb{K})$ is definable
It is not hard to see that given a $T$-convex subring $\mathcal{O}\subseteq\mathbb{E}\models T$, there is $\mathbb{K}\preceq\mathbb{E}$ maximal in $\{\mathbb{K}^{\prime}\preceq\mathbb{E}:\mathbb{K}\subseteq\mathcal{O}\}$.
Fact 7 (van den Dries [1], van den Dries and Lewenberg [3]).
If $(\mathbb{E},\mathcal{O})\models T_{\mathrm{convex}}^{-}$, then the following are equivalent for $\mathbb{K}\preceq\mathbb{E}$:
  1. (1)
    $\mathbb{K}$ is maximal in $\{\mathbb{K}^{\prime}\preceq\mathbb{E}:\mathbb{K}\subseteq\mathcal{O}\}$;
  2. (2)
    $\mathbb{K}\preceq_{\mathrm{tame}}\mathbb{E}$ and $\operatorname{\mathrm{CH}}(\mathbb{K})=\mathcal{O}$;
  3. (3)
    $\mathfrak{o}\cap\mathbb{K}=\{0\}$ and $\operatorname{\mathrm{st}}_{\mathbb{K}}:\mathcal{O}\to\mathbb{K}$ induces an isomorphism between the induced structure on the imaginary sort $\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$ and $\mathbb{K}$.
Proof.
See [3, 2.12] for the equivalence of (1) and (2). The equivalence of (2) and (3) is instead in [1, 1.13]. ∎
Definition 8.
We call $\mathbb{K}\prec\mathbb{E}$ an elementary residue-section for $(\mathbb{E},\mathcal{O})$ when the equivalent conditions of Fact 7 hold.
Fact 9 (van den Dries and Lewenberg (5.3) in [3]).
Suppose $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}\langle x\rangle,\mathcal{O}^{\prime})\models T_{\mathrm{convex}}$, then $\dim_{\operatorname{\mathrm{dcl}}}(\operatorname{\mathbf{res}}(\mathbb{E}\langle x\rangle,\mathcal{O}^{\prime})/\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O}))\leq 1$.
2.3. Cuts and $T$-convex valuations
Let $(\mathbb{U},\mathcal{O})\models T_{\mathrm{convex}}$ and $\mathbb{E}\prec\mathbb{U}\models T$.
Definition 10.
We call an element $x\in\mathbb{U}\setminus\mathbb{E}$:
  1. (1)
    weakly $\mathcal{O}$-immediate (or $\mathcal{O}$-wim) over $\mathbb{E}$ if $\operatorname{\mathbf{v}}_{\mathcal{O}}(x-\mathbb{E})$ has no maximum;
  2. (2)
    $\mathcal{O}$-residual over $\mathbb{E}$ if $x\in\mathcal{O}$ and $\operatorname{\mathbf{res}}_{\mathcal{O}}(x)\notin\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$;
  3. (3)
    $\mathcal{O}$-purely valuational over $\mathbb{E}$ if $\mathbb{E}\langle x\rangle\setminus\mathbb{E}=\mathbb{E}+M$ for some $M\subseteq\mathbb{E}\langle x\rangle$ such that $\operatorname{\mathbf{v}}_{\mathcal{O}}(M)\cap\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E})=\emptyset$.
Furthermore we will say that $x$ is weakly $\mathcal{O}$-immediately generated (or $\mathcal{O}$-wim generated) over $\mathbb{E}$ if $\mathbb{E}\langle x\rangle=\mathbb{E}\langle y\rangle$ for some weakly $\mathcal{O}$-immediate $x$. Similarly we will say that $x$ is $\mathcal{O}$-residually generated over $\mathbb{E}$ if $\mathbb{E}\langle x\rangle=\mathbb{E}\langle y\rangle$ for some residual $y$.
Remark 11.
An element $x\in\mathbb{U}\setminus\mathbb{E}$ is weakly immediate over $\mathbb{E}$, if and only if for every $c\in\mathbb{E}$, $\operatorname{\mathbf{rv}}_{\mathcal{O}}(x-c)\in\operatorname{\mathbf{rv}}_{\mathcal{O}}(\mathbb{E})$.
Fact 12 (see Thm. 3.10 in [6]).
For every $x\in\mathbb{U}\setminus\mathbb{E}$ only one of the following holds:
  1. (1)
    $x$ is $\mathcal{O}$-wim-generated;
  2. (2)
    $x$ is $\mathcal{O}$-residually generated;
  3. (3)
    $x$ is $\mathcal{O}$-purely valuational.
In particular if $x$ is $\mathcal{O}$-residually generated over $\mathbb{E}$, then $\mathbb{E}\langle x\rangle$ does not contain elements that are $\mathcal{O}$-wim over $\mathbb{E}$.
Recall from [8, Sec. 1] that if $\mathbb{E}$ is an o-minimal expansion of a field, then a power-function of $\mathbb{E}$ is a definable endomorphism of $(\mathbb{E}^{>0},\cdot)$; by o-minimality a power-function $\theta:\mathbb{E}^{>0}\to\mathbb{E}^{>0}$ is always monotone and differentiable, its exponent $\lambda$ is then defined $\theta^{\prime}(1)$ and $\theta(x)$ is usually written as $x^{\lambda}$.
Also recall that the exponents of (the power-functions of) $\mathbb{E}$ always form a subfield of $\mathbb{E}$ and recall from [8, Thm. 3.6] that for every complete o-minimal theory $T$ exactly one of the following holds:
  • $T$ is power bounded: for every model $\mathbb{E}\models T$, all power functions of $\mathbb{E}$ are $\emptyset$-definable and all $\mathbb{E}$-definable total unary functions are eventually bounded by some power-function;
  • $T$ is exponential: for every model $\mathbb{E}\models T$, there is a $T$-definable ordered group isomorphism $\exp:(\mathbb{E},+,<)\to(\mathbb{E}^{>0},\cdot,<)$ such that $\exp^{\prime}(0)=0$.
When $T$ is power-bounded a stronger version of Fact 12 holds: this is known as the residue-valuation property or rv-property of power-bounded o-minimal theories.
Fact 13 (Tyne [10], van den Dries and Speissegger [5]).
If $T$ is power bounded with field of exponents $\Lambda$, then for every $x\in\mathbb{U}\setminus\mathbb{E}$, only one of the following holds
  1. (1)
    $x$ is $\mathcal{O}$-wim;
  2. (2)
    there are $c,d\in\mathbb{E}^{\neq 0}$ such that $\operatorname{\mathbf{res}}(d(x-c))\notin\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$;
  3. (3)
    there is $c\in\mathbb{E}$ such that $\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E}\langle x\rangle)=\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E})+\Lambda\cdot\operatorname{\mathbf{v}}_{\mathcal{O}}(x-c)$.
3. Main results
3.1. Res-constructible extensions
We recall the definition of res-constructible extension from the introduction and give some of its basic properties.
Definition 14.
Let $\mathbb{E}\preceq\mathbb{E}_{*}\models T$ and let $\mathcal{O}_{*}$ be a non-trivial $T$-convex subring of $\mathbb{E}_{*}$. A possibly infinite tuple $\overline{y}:=(y_{i}:i\in I)$ in $\mathcal{O}_{*}$ will be called a $\mathcal{O}_{*}$-res-construction over $\mathbb{E}$ if $\big(\operatorname{\mathbf{res}}_{\mathcal{O}_{*}}(y_{i}):i\in I\big)$ is $\operatorname{\mathrm{dcl}}_{T}$-independent over $\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$. We then call the extension $\mathbb{E}\langle y_{i}:i\in I\rangle$ of $\mathbb{E}$ $\mathcal{O}_{*}$-res-constructible over $\mathbb{E}$ (by the res-construction $(y_{i}:i\in I)$). When $\mathcal{O}_{*}$ is clear from the context we will just say res-construction and res-constructible. We call an extension $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})\models T_{\mathrm{convex}}$ residually constructible (res-constructible for short) when $\mathbb{E}_{*}$ is $\mathcal{O}_{*}$-res constructible over $\mathbb{E}$.
Remark 15.
Notice that if $\overline{y}:=(y_{i}:i\in I)$ is a $\mathcal{O}_{*}$-res-construction over $\mathbb{E}$, then it is in particular $\operatorname{\mathrm{dcl}}_{T}$-independent over $\mathbb{E}$. This is because by Fact 9, for all finite $F\subseteq I$, $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}\langle y_{i}:i\in F\rangle)=\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})\langle\operatorname{\mathbf{res}}(y_{i}):i\in F\rangle$.
Remark 16.
Res-constructible extension are transitive (i.e. closed under composition): if $(\mathbb{E}_{0},\mathcal{O}_{0})\preceq(\mathbb{E}_{1},\mathcal{O}_{1})$ and $(\mathbb{E}_{1},\mathcal{O}_{1})\preceq(\mathbb{E}_{2},\mathcal{O}_{2})$ are res-constructible, then $(\mathbb{E}_{0},\mathcal{O}_{0})\preceq(\mathbb{E}_{2},\mathcal{O}_{2})$ is res-constructible as well.
Remark 17.
In the same setting of Definition 14, let $\lambda$ be an ordinal, and $(y_{i}:i<\lambda)$ a $\lambda$-indexed sequence in $\mathcal{O}_{*}$. Then the following are equivalent:
  1. (1)
    $(y_{i}:i<\lambda)$ is a $\mathcal{O}_{*}$-res-construction over $\mathbb{E}$;
  2. (2)
    for all $i<\lambda$, we have $\operatorname{\mathbf{res}}_{\mathcal{O}_{*}}(y_{i})\in\operatorname{\mathbf{res}}(\mathbb{E}\langle y_{i}:j<i+1\rangle)\setminus\operatorname{\mathbf{res}}_{\mathcal{O}_{*}}(\mathbb{E}\langle y_{j}:j<i\rangle)$.
In particular, by Fact 9, for an extension $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})$ to be res-constructible, it is enough that for some $\operatorname{\mathrm{dcl}}_{T}$-basis $B$ of $\mathbb{E}_{*}$ over $\mathbb{E}$, there is an ordinal indexing $(b_{i}:i<\alpha)$ of $B$ such that for all $i<\alpha$, one has $\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{j}:j<i\rangle)\neq\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{j}:j<i+1\rangle)$.
Lemma 18.
If $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})$ is res-constructible by the ordinal indexed res-construction $(y_{i}:i<\lambda)$, and $\mathbb{K}$ is an elementary residue section of $(\mathbb{E},\mathcal{O})$, then $\mathbb{K}\langle y_{i}:i<\lambda\rangle$ is an elementary residue section of $(\mathbb{E}_{*},\mathcal{O}_{*})$.
Proof.
Notice that by Fact 7 (3) it suffices to show that $\operatorname{\mathbf{res}}(\mathbb{K}\langle y_{i}:i<\lambda\rangle)=\operatorname{\mathbf{res}}(\mathbb{E}_{*},\mathcal{O}_{*})$. Proceed by induction on $\lambda$. Let $\lambda=\beta+1$ and suppose the thesis holds for $\beta$, so assume that $\mathbb{K}\langle y_{i}:i<\lambda\rangle$ is an elementary a residue section of $(\mathbb{E}\langle y_{i}:i<\beta\rangle,\mathcal{O}_{*}\cap\mathbb{E}\langle y_{i}:i<\beta\rangle)$. Since by Fact 9 we must then have $\operatorname{\mathbf{res}}(\mathbb{E}\langle y_{i}:j<\lambda\rangle)=\operatorname{\mathbf{res}}(\mathbb{E}\langle y_{i}:j<\beta\rangle)\langle\operatorname{\mathbf{res}}(y_{\beta})\rangle$, we deduce by Fact 7 that $\operatorname{\mathbf{res}}(\mathbb{K}\langle y_{i}:i<\lambda\rangle)=\operatorname{\mathbf{res}}(\mathbb{E}\langle y_{i}:j<\lambda\rangle)$.
Suppose instead that $\lambda$ is a limit ordinal and the thesis holds for each $\beta<\lambda$. If $x\in\mathbb{E}_{*}$, then since $\lambda$ is a limit ordinal, there is $\beta<\lambda$ such that $x\in\mathbb{E}\langle y_{i}:i<\beta\rangle$ and hence by inductive hypothesis $\operatorname{\mathbf{res}}(x)\in\operatorname{\mathbf{res}}(\mathbb{K}\langle y_{i}:i<\beta\rangle)\subseteq\operatorname{\mathbf{res}}(\mathbb{K}\langle y_{i}:i<\lambda\rangle)$. ∎
The isomorphism type of a res-constructible extension only depends on the isomorphism type of the residue field extension $\operatorname{\mathbf{res}}(\mathbb{E}_{*},\mathcal{O}_{*})\succeq\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$.
Lemma 19.
If $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{1},\mathcal{O}_{1})$ and $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{2},\mathcal{O}_{2})$ are res-constructible extensions and $\varphi:\operatorname{\mathbf{res}}(\mathbb{E}_{1},\mathcal{O}_{1})\to\operatorname{\mathbf{res}}(\mathbb{E}_{2},\mathcal{O}_{2})$ is an isomorphism of models of $T$ over $\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$, then there is an isomorphism $\psi:(\mathbb{E}_{1},\mathcal{O}_{1})\to(\mathbb{E}_{2},\mathcal{O}_{2})$ over $\mathbb{E}$ such that $\operatorname{\mathbf{res}}(\psi)=\varphi$.
Proof.
Let $\mathbb{K}$ be an elementary residue-section of $(\mathbb{E},\mathcal{O})$ and $(y_{i}:i<\lambda)$, $(z_{i}:i<\mu)$ be res-construction for $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$. Thus $\mathbb{E}=\mathbb{K}\langle\overline{d}\rangle$ where $\overline{d}$ is a tuple such that $\operatorname{\mathrm{tp}}(\overline{d}/\mathbb{K})$ is definable. By Lemma 18, $\mathbb{K}_{1}:=\mathbb{K}\langle y_{i}:i<\lambda\rangle$ and $\mathbb{K}_{2}:=\mathbb{K}\langle z_{i}:i<\mu\rangle$ are elementary residue sections for $(\mathbb{E}_{1},\mathcal{O}_{1})$ and $(\mathbb{E}_{2},\mathcal{O}_{2})$ respectively and moreover $\mathbb{E}_{1}=\mathbb{K}_{1}\langle\overline{d}\rangle$ and $\mathbb{E}_{2}=\mathbb{K}_{2}\langle\overline{d}\rangle$. Now observe that $\varphi$ induces an isomorphism $h:\mathbb{K}_{1}\to\mathbb{K}_{2}$ over $\mathbb{K}$. Finally notice that for both $i=1$ and $i=2$, $\operatorname{\mathrm{tp}}(\overline{d}/\mathbb{K}_{i})$ is the unique definable extension of $\operatorname{\mathrm{tp}}(\overline{d}/\mathbb{K})$ to $\mathbb{K}_{i}$, in particular $h_{*}\operatorname{\mathrm{tp}}(\overline{d}/\mathbb{K}_{1})=\operatorname{\mathrm{tp}}(\overline{d}/\mathbb{K}_{2})$. Thus $h$ extends to the sought isomorphism $\psi:(\mathbb{E}_{1},\mathcal{O}_{1})\to(\mathbb{E}_{2},\mathcal{O}_{2})$ over $\mathbb{E}$. ∎
Res-constructible extension only realize residually generated elements.
Lemma 20.
If $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})$ is a res-constructible extension then every $x\in\mathbb{E}_{*}\setminus\mathbb{E}$ is $\mathcal{O}_{*}$-residually generated over $\mathbb{E}$.
Proof.
Let $(r_{i}:i<\alpha)$ be an ordinal-indexed res-construction for $\mathbb{E}_{*}$ over $\mathbb{E}$. Let $j$ be minimal such that $\operatorname{\mathrm{tp}}(x/\mathbb{E})$ is realized in $\mathbb{E}\langle r_{i}:i<j+1\rangle$. Thus there is $x^{\prime}\in\mathbb{E}\langle r_{i}:i<j+1\rangle\setminus\mathbb{E}\langle r_{i}:i<j\rangle$ such that $\operatorname{\mathrm{tp}}(x^{\prime}/\mathbb{E})=\operatorname{\mathrm{tp}}(x/\mathbb{E})$. Now let $F\subseteq j$ such that $x^{\prime}\in\mathbb{E}\langle r_{i}:i\in F\cup\{j\}\rangle$. By Fact 9 we then have
$\operatorname{\mathbf{res}}(\mathbb{E}\langle x^{\prime},r_{i}:i\in F\rangle)=\operatorname{\mathbf{res}}(\mathbb{E}\langle r_{i}:i\in F\rangle)\langle\operatorname{\mathbf{res}}(r_{j})\rangle.$
In particular $\dim_{\operatorname{\mathrm{dcl}}}(\operatorname{\mathbf{res}}(\mathbb{E}\langle x^{\prime},r_{i}:i\in F\rangle)/\operatorname{\mathbf{res}}(\mathbb{E}))=|F|+1$. Suppose toward contradiction that $x^{\prime}$ was not residually generated over $\mathbb{E}$, i.e. that $\operatorname{\mathbf{res}}(\mathbb{E}\langle x^{\prime}\rangle)=\operatorname{\mathbf{res}}(\mathbb{E})$, then, again by Fact 9, we would have $\dim_{\operatorname{\mathrm{dcl}}}(\operatorname{\mathbf{res}}(\mathbb{E}\langle x^{\prime},r_{i}:i<F\rangle/\operatorname{\mathbf{res}}(\mathbb{E}))\leq|F|$, contradiction. Thus $x^{\prime}$ is residually generated over $\mathbb{E}$ and so also $x$ is, because $\operatorname{\mathrm{tp}}(x/\mathbb{E})=\operatorname{\mathrm{tp}}(x^{\prime}/\mathbb{E})$. ∎
Corollary 21.
If $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{*},\mathcal{O}_{*})$ is a res-constructible extension and $\overline{y}$ is a finite tuple, then $\mathbb{E}_{*}$ is $\mathcal{O}_{*}$-res-constructible over $\mathbb{E}\langle\overline{y}\rangle$ and $\mathbb{E}\langle\overline{y}\rangle$ is $\mathcal{O}_{*}$-res-constructible over $\mathbb{E}$.
Proof.
It suffices to prove the case in which $\overline{y}=(y_{0})$ is a single element of $\mathbb{E}_{*}$. If $y_{0}\notin\mathbb{E}$, then by Lemma 20, $\mathbb{E}\langle y_{0}\rangle$ is res-constructible over $\mathbb{E}$, say by some $r\in\mathbb{E}\langle y_{0}\rangle$, such that $\operatorname{\mathbf{res}}(r)\notin\operatorname{\mathbf{res}}(\mathbb{E})$. Let $\mathbb{K}$ be an elementary residue section of $(\mathbb{E},\mathcal{O})$ and let $\mathbb{K}_{*}$ an extension of $\mathbb{K}$ to an elementary residue section of $\mathbb{E}_{*}$ such that $\mathbb{E}_{*}=\operatorname{\mathrm{dcl}}(\mathbb{E}\cup\mathbb{K}_{*})$, which exists by Lemma 18. Now if $\overline{s}$ is a $\operatorname{\mathrm{dcl}}_{T}$-basis of $\mathbb{K}_{*}$ over $\mathbb{K}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(r)\rangle$, then $\overline{s}$ is a $\mathcal{O}_{*}$-res-construction for $\mathbb{E}_{*}$ over $\mathbb{E}\langle r\rangle=\mathbb{E}\langle y_{0}\rangle$. In fact $\operatorname{\mathbf{res}}(r,\overline{s})=\operatorname{\mathbf{res}}(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(r),\overline{s})$ is $\operatorname{\mathrm{dcl}}$-independent over $\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$; and since then $r\notin\mathbb{E}\langle\overline{s}\rangle$, we also have by the exchange property that $\mathbb{E}\langle r,\overline{s}\rangle=\mathbb{E}_{*}$. ∎
Lemma 22.
Let $(\mathbb{E},\mathcal{O})\models T_{\mathrm{convex}}$, and let $x$ be an element of an elementary extension $(\mathbb{E}_{*},\mathcal{O}_{*})$, such that $x$ is not residually generated over $(\mathbb{E},\mathcal{O})$. If $\operatorname{\mathbf{v}}(\mathbb{E},\mathcal{O})$ does not contain any uncountable ordinal, then there is a countable $S\subseteq\mathbb{E}$, such that $\operatorname{\mathrm{tp}}_{L_{\mathrm{convex}}}(x/S)\vdash\operatorname{\mathrm{tp}}_{L_{\mathrm{convex}}}(x/\mathbb{E})$.
Proof.
[…]
In Section 4 we will need the following two easy facts about res-constructible extensions with respect to different $T$-convex valuations.
Lemma 23.
Let $\mathbb{E}\preceq\mathbb{U}\models T$ and let $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ be $T$-convex valuation subrings of $\mathbb{U}$. If $\overline{x}$ is a $\mathcal{O}$-res-construction over $\mathbb{E}$ then $\overline{x}$ is also an $\mathcal{O}^{\prime}$-res-construction over $\mathbb{E}$ and $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\overline{x})$ is a $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathcal{O})$-res construction over $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$.
Proof.
Notice that $V:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathcal{O})$ is a $T$-convex valuation subring of $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{U})$ and that $\operatorname{\mathbf{res}}_{\mathcal{O}}$ factors through $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}$ as $\operatorname{\mathbf{res}}_{\mathcal{O}}=\pi\circ\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}$ where $\pi$ is in fact $\operatorname{\mathbf{res}}_{V}$. Now suppose that $\overline{x}:=(x_{i}:i\in I)$ is such that $(\operatorname{\mathbf{res}}_{\mathcal{O}}(x_{i}):i\in I)$ is $\operatorname{\mathrm{dcl}}$-independent over $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$. It then follows from Remark 15 that $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\overline{x})$ is independent over $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$ because $\operatorname{\mathbf{res}}_{\mathcal{O}}(\overline{x})=\pi(\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\overline{x}))$ is $\operatorname{\mathrm{dcl}}_{T}$-independent. ∎
Corollary 24.
Let $\mathbb{E}\prec\mathbb{U}\models T$, $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ be $T$-convex valuation subrings of $\mathbb{U}$, and $y\in\mathbb{U}\setminus\mathbb{E}$. If $y$ is $\mathcal{O}$-residually generated, then it is $\mathcal{O}^{\prime}$-residually generated.
3.2. Right factors of res-constructible extensions
This subsection is devoted to the proof of Theorem A, which will result as a wrap-up of Theorems 35 and 36. Throughout this subsection we fix a res-constructible extension $(\mathbb{E}_{*},\mathcal{O}_{*})\succeq(\mathbb{E},\mathcal{O})$ of model of $T_{\mathrm{convex}}$ and let $\overline{r}$ be a $\mathcal{O}_{*}$-res-construction for $\mathbb{E}_{*}$ over $\mathbb{E}$. We also let $\mathbb{E}_{1}$ be an intermediate extension $\mathbb{E}\preceq\mathbb{E}_{1}\preceq\mathbb{E}_{*}$.
Definition 25.
Let $\mathbb{K}\preceq_{\mathrm{tame}}\mathbb{E}$ be a residue section for $(\mathbb{E},\mathcal{O})$ and $\mathbb{K}_{*}:=\mathbb{K}\langle\overline{r}\rangle$, thus $\mathbb{K}_{*}$ is an elementary residue section for $(\mathbb{E}_{*},\mathcal{O}_{*})$ by Lemma 18. We will say that a $\mathcal{O}_{*}$-res-construction $\overline{x}$ in $\mathbb{E}_{1}$ over $\mathbb{E}$ is
  1. (1)
    $\mathbb{K}_{*}$-orthogonal in $\mathbb{E}_{1}$ if $\operatorname{\mathrm{tp}}_{L_{\mathrm{convex}}}(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x})/\mathbb{E}\langle\overline{x}\rangle)\vdash\operatorname{\mathrm{tp}}_{L_{\mathrm{convex}}}(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x})/\mathbb{E}_{1})$;
  2. (2)
    $\mathbb{K}_{*}$-strong in $\mathbb{E}_{1}$ (notation $\overline{x}\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$) if it is $\mathbb{K}_{*}$-orthogonal in $\mathbb{E}_{1}$, and $\mathbb{E}\langle\overline{x}\rangle=\mathbb{E}_{1}\cap\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}^{*}}(\overline{x})\rangle$.
Lemma 26.
Let $\overline{x}\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$, and suppose $\psi:(\mathbb{E}_{*},\mathcal{O}_{*})\to(\mathbb{E}_{*},\mathcal{O}_{*})$ is an elementary self-embedding fixing $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x})$ and $\mathbb{E}$ elementwise. Then $\psi(\overline{x})\triangleleft_{\mathbb{K}_{*}}\psi(\mathbb{E}_{1})$.
Proof.
Since $\overline{x}\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$, then $\overline{x}\subset\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x})\rangle$. Thus, $\psi$ also fixes $\overline{x}$ elementwise. This implies that that $\psi(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}))=\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\psi(\overline{x}))$. Then, by applying the elementarity of $\psi$ and the fact that $x$ is $\mathbb{K}_{*}$-orthogonal in $\mathbb{E}_{1}$, we deduce that
$\operatorname{\mathrm{tp}}\big(\psi(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}))/\psi(\mathbb{E})\langle\psi(\overline{x})\rangle\big)\vdash\operatorname{\mathrm{tp}}\big(\psi(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}))/\psi(\mathbb{E}_{1})\big).$
Applying $\psi(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}))=\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\psi(\overline{x}))$ yields that $\psi(\overline{x})$ is $\mathbb{K}_{*}$-orthogonal in $\psi(\mathbb{E}_{1})$. Finally observe that $\mathbb{E}\langle\psi(\overline{x})\rangle=\psi(\mathbb{E}\langle\overline{x}\rangle)=\psi(\mathbb{E}_{1})\cap\psi(\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x})\rangle)=\psi(\mathbb{E}_{1})\cap\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\psi(\overline{x}))\rangle$. ∎
Lemma 27.
Let $\overline{x}$ be a $\mathbb{K}_{*}$-orthogonal $\mathcal{O}_{*}$-res-construction in $\mathbb{E}_{1}$ and $\overline{y}$ a $\mathcal{O}_{*}$-res-construction in $\mathbb{E}_{1}$ over $\mathbb{E}\langle\overline{x}\rangle$. If $z\in\mathbb{E}_{1}$ is $\mathcal{O}_{*}$-residually generated over $\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}),\overline{y}\rangle$ then it is $\mathcal{O}_{*}$-residually generated over $\mathbb{E}\langle\overline{x},\overline{y}\rangle$.
Proof.
Notice that $\operatorname{\mathbf{res}}(\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}),\overline{y}\rangle)=\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{x},\overline{y}\rangle)\subseteq\operatorname{\mathbf{res}}(\mathbb{E}_{1})$. Suppose that $z$ is residually generated over $\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}),\overline{y}\rangle$. Then there is a finite subtuple $\overline{s}$ of $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x})$ and a $\mathbb{E}\langle\overline{y}\rangle$-definable function $f$ such that $\operatorname{\mathbf{res}}(f(\overline{s},z))\notin\operatorname{\mathbf{res}}(\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}),\overline{y}\rangle)$. However $\operatorname{\mathbf{res}}(f(\overline{s},z))=\operatorname{\mathbf{res}}(a)$ for some $a\in\mathbb{E}_{1}$ and by $\mathbb{K}_{*}$-orthogonality of $\overline{x}$ in $\mathbb{E}_{1}$, we have $\operatorname{\mathrm{tp}}_{L_{\mathrm{convex}}}(s/\mathbb{E}\langle\overline{x}\rangle)\vdash\operatorname{\mathbf{res}}(f(\overline{s},z))=\operatorname{\mathbf{res}}(a)$. But then there is a $L_{\mathrm{convex}}$-formula $\varphi(\overline{s})$ with parameters from $\mathbb{E}\langle\overline{x}\rangle$ such that $\varphi(\overline{s})\vdash\operatorname{\mathbf{res}}(f(\overline{s},z))=\operatorname{\mathbf{res}}(a)$. Taking $\overline{s}^{\prime}$ in $\mathbb{E}\langle\overline{x}\rangle$ such that $\big(\mathbb{E}\langle\overline{x}\rangle,\mathcal{O}_{*}\cap\mathbb{E}\langle\overline{x}\rangle\big)\models\varphi(\overline{s}^{\prime})$ we thus get $\operatorname{\mathbf{res}}(f(\overline{s}^{\prime},z))=\operatorname{\mathbf{res}}(a)=\operatorname{\mathbf{res}}(f(\overline{s},z))$, whence $z$ is $\mathcal{O}_{*}$-residually generated over $\mathbb{E}\langle\overline{x},\overline{y}\rangle$. ∎
Lemma 28.
Let $\overline{x}\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$ and let $\overline{y}$ be a finite res-construction in $\mathbb{E}_{1}$ over $\mathbb{E}\langle\overline{x}\rangle$. Then every $z\in\mathbb{E}_{1}\setminus\mathbb{E}\langle\overline{x},\overline{y}\rangle$ is $\mathcal{O}_{*}$-residually generated over $\mathbb{E}\langle\overline{x},\overline{y}\rangle$.
Proof.
Let $\overline{r}^{\prime}$ be a completion of $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x},\overline{y})$ to a $\operatorname{\mathrm{dcl}}_{T}$-basis of $\mathbb{K}_{*}$. Set $\psi(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{y})):=\overline{y}$, $\psi(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x})):=\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x})$ and $\psi_{\mathbb{K}_{*}}(\overline{r}^{\prime})=\overline{r}^{\prime}$. Notice that $\psi$ extends to an elementary self-embedding of $(\mathbb{E}_{*},\mathcal{O}_{*})$ over $\mathbb{E}$ and that it fixes $\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}),\overline{r}^{\prime}\rangle$ which has finite codimension in $\mathbb{E}_{*}$, therefore $\psi$ is an automorphism. By Lemma 26, we have $\psi^{-1}(\overline{x})\triangleleft_{\psi^{-1}(\mathbb{K}_{*})}\psi^{-1}(\mathbb{E}_{1})$. Moreover $\psi^{-1}(z)\in\psi^{-1}(\mathbb{E}_{1})$ must be residual over $\psi^{-1}(\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}),\overline{y}\rangle)=\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}),\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{y})\rangle$ whence by Lemma 27 it is $\mathcal{O}_{*}$-residually generated over $\psi^{-1}(\mathbb{E}\langle\overline{x},\overline{y}\rangle)$. Applying $\psi$ finally yields that $z$ is $\mathcal{O}_{*}$-residual over $\mathbb{E}\langle\overline{x},\overline{y}\rangle$. ∎
As a Corollary of Lemma 28, we get that if $\overline{x}\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$, then for all $\mathbb{E}_{0}$ such that $\mathbb{E}\langle\overline{x}\rangle\preceq\mathbb{E}_{0}\preceq\mathbb{E}_{1}$ and $\dim_{\operatorname{\mathrm{dcl}}}(\mathbb{E}_{0}/\mathbb{E}\langle\overline{x}\rangle)\leq\aleph_{0}$, we have that $\mathbb{E}_{0}$ is $\mathcal{O}_{*}$-res-constructible over $\mathbb{E}\langle\overline{x}\rangle$. We give a more precise statement of this.
Corollary 29.
Suppose that $\overline{x}\trianglelefteq_{\mathbb{K}_{*}}\mathbb{E}_{1}$ and $\overline{y}:=(y_{i}:i<\alpha)$ in $\mathbb{E}_{1}$ is $\operatorname{\mathrm{dcl}}_{T}$-independent over $\mathbb{E}\langle\overline{x}\rangle$ with $\alpha\leq\omega$. Then there is a $\mathcal{O}_{*}$-res-construction $\overline{z}:=(z_{i}:i<\alpha)$ in $\mathbb{E}\langle\overline{x},\overline{y}\rangle$ over $\mathbb{E}\langle\overline{x}\rangle$, such that for all $n\leq\alpha$, $\mathbb{E}\langle\overline{x},x_{i}:i<n\rangle=\mathbb{E}\langle\overline{x},y_{i}:i<n\rangle$.
Proof.
By induction, using Lemma 28, we can build a $\mathcal{O}_{*}$-res-construction $(z_{i}:i<\alpha)$ over $\mathbb{E}\langle\overline{x}\rangle$ such that for all $n<\alpha$, we have $\mathbb{E}\langle\overline{x},z_{i}:i<n\rangle=\mathbb{E}\langle\overline{x},y_{i}:i<n\rangle$. ∎
$\mathbb{K}_{*}$-strong res-constructions are closed under directed unions.
Lemma 30.
Let $\overline{x}:=(x_{i}:i\in I)$ be family of elements in $\mathcal{O}_{*}\cap\mathbb{E}_{1}$. If for all finite $F\subseteq I$, $\overline{x}_{F}:=(x_{i}:i\in F)\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$, then $\overline{x}\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$.
Proof.
The fact that $\overline{x}$ is a $\mathcal{O}_{*}$-res/construction follows from the finite character of $\operatorname{\mathrm{dcl}}_{T}$-independence. Similarly, that $\overline{x}$ is $\mathbb{K}_{*}$-orthogonal follows from the fact that every formula in $\operatorname{\mathrm{tp}}(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x})/\mathbb{E}_{1})$ is in fact in some $\operatorname{\mathrm{tp}}(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}_{F})/\mathbb{E}_{1})$ for some finite $F\subseteq I$. Finally $\mathbb{E}\langle\overline{x}\rangle=\bigcup_{F}\mathbb{E}\langle\overline{x}_{F}\rangle=\mathbb{E}_{1}\cap\bigcup_{F}\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x})\rangle$, where in the unions $F$ ranges among the finite subsets of $I$. ∎
Lemma 31.
Suppose that $\operatorname{\mathbf{v}}_{\mathcal{O}_{*}}(\mathbb{E}_{1})$ contains no uncountable well ordered set and $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}_{1})=\mathbb{K}_{*}$. Let $\mathbb{E}\preceq\mathbb{E}_{0}\preceq\mathbb{E}_{1}$ and $\overline{y}$ be a finite $\mathcal{O}_{*}$-res-construction in $\mathbb{E}_{1}$ over $\mathbb{E}_{0}$. Then there is a countable tuple $\overline{z}$ in $\mathbb{E}_{1}$ such that $\operatorname{\mathrm{tp}}_{L_{\mathrm{convex}}}(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{y})/\mathbb{E}_{0}\langle\overline{z}\rangle)\vdash\operatorname{\mathrm{tp}}_{L_{\mathrm{convex}}}(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{y})/\mathbb{E}_{1})$.
Proof.
Notice that we can reduce to the case in which $y$ is a single residual element. Let $s=\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(y)$. We can assume that $s\notin\mathbb{E}_{1}$ for otherwise the thesis is trivial. Since $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}_{1})=\mathbb{K}_{*}$, $s$ is not residually generated over $\mathbb{E}_{1}$, so the thesis follows from the fact that $\operatorname{\mathbf{v}}_{\mathcal{O}_{*}}(\mathbb{E}_{1})$ contains no uncountable well ordered subset and from Lemma 22. ∎
Proposition 32.
Let $\overline{x}\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$ and let $B$ be a $\operatorname{\mathrm{dcl}}_{T}$-basis of $\mathbb{E}_{1}$ over $\mathbb{E}\langle\overline{x}\rangle$. Suppose that $\operatorname{\mathbf{v}}_{\mathcal{O}_{*}}(\mathbb{E}_{1})$ contains no uncountable ordinal and that $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}_{1})=\mathbb{K}_{*}$. If $B$ is infinite, then every $b_{0}\in B$ extends to a countable sequence $(b_{i}:i<\omega)$ in $B$ such that there is $\overline{x}^{\prime}:=(x^{\prime}_{i}:i<\omega)$ with $(\overline{x},\overline{x}^{\prime})\trianglelefteq_{\mathbb{K}_{*}}\mathbb{E}_{1}$ and $\mathbb{E}\langle\overline{x},x_{i}^{\prime}:i<n\rangle=\mathbb{E}\langle\overline{x},b_{i}:i<n\rangle$ for all $n\leq\omega$.
Proof.
Consider functions $f,g:[B]^{<\aleph_{0}}\to[B]^{<\aleph_{1}}$ defined as follows.
  • given a finite subset $S\subseteq B$, let $S^{\prime}$ be a $\operatorname{\mathrm{dcl}}_{T}$-basis of $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}\langle\overline{x},S\rangle)$ over $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}\langle\overline{x}\rangle)$. Set $f(S)$ to be a countable subset of $B$ such that
    $\operatorname{\mathrm{tp}}_{L_{\mathrm{convex}}}(S^{\prime}/\mathbb{E}\langle\overline{x},f(S)\rangle)\vdash\operatorname{\mathrm{tp}}_{L_{\mathrm{convex}}}(S^{\prime}/\mathbb{E}_{1});$
  • given a finite subset $S\subseteq B$, choose a finite $R\subseteq\mathbb{K}_{*}$ such that $S\subseteq\mathbb{E}\langle R\rangle$, then set $g(S)$ to be such that $R\subseteq\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}\langle\overline{x},g(S)\rangle)$ and $\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}),R\rangle\cap\mathbb{E}_{1}\subseteq\mathbb{E}\langle\overline{x},g(S)\rangle$ (notice that one can choose a finite such $g(S)$).
By a standard diagonalization argument, given $b_{0}\in B$ there is a countable subset $\{b_{i}:i<\omega\}$ of $B$ closed under $f$ and $g$. Let $\overline{x}^{\prime}$ be a res-construction of $\mathbb{E}\langle\overline{x},b_{i}:i<\omega\rangle$ over $\mathbb{E}\langle\overline{x}\rangle$. To see it is $\mathbb{K}_{*}$-orthogonal notice that $\operatorname{\mathrm{tp}}(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}^{\prime})/\mathbb{E}\langle\overline{x},\overline{x}^{\prime}\rangle)\vdash\operatorname{\mathrm{tp}}(\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}^{\prime})/\mathbb{E}_{1})$ by compactness and closure under $f$. To see it is $\mathbb{K}_{*}$-strong, notice that $\mathbb{E}\langle\overline{x},\overline{x}^{\prime}\rangle\subseteq\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}),\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}^{\prime})\rangle$ and $\mathbb{E}\langle\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}),\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\overline{x}^{\prime})\rangle\cap\mathbb{E}_{1}\subseteq\mathbb{E}\langle\overline{x},\overline{x}^{\prime}\rangle$ by closure under $g$. ∎
An immediate Corollary of Proposition 32, is the following, which amounts to the first part of our main theorem, but with extra assumption on $\mathbb{E}_{1}$ that $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}_{1})=\mathbb{K}_{*}$.
Corollary 33.
Suppose that $\overline{x}\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$, $\operatorname{\mathbf{v}}_{\mathcal{O}_{*}}(\mathbb{E}_{1})$ contains no uncountable ordinal and $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}_{1})=\mathbb{K}_{*}$. Then for every $\operatorname{\mathrm{dcl}}_{T}$-basis $B$ of $\mathbb{E}_{1}$ over $\mathbb{E}\langle\overline{x}\rangle$, there is an ordinal-indexing $(b_{i}:i<\lambda)$ of $B$, such that for all $i<\lambda$, $\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{x},b_{j}:j<i\rangle)\neq\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{x},b_{j}:j\leq i\rangle)$.
Proof.
Using Proposition 32 and Lemma 30 one can find some ordinal $\alpha$, $\overline{x}^{\prime}:=(x_{i}^{\prime}:i<\alpha)$ in $\mathbb{E}_{1}$ and $(b_{i}:i<\alpha)$ in $B$, such that $\{b_{i}:i<\alpha\}$ is cofinite in $B$, $(\overline{x},\overline{x}^{\prime})\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$, and for all $i\leq\alpha$, $\mathbb{E}\langle\overline{x},x_{j}^{\prime}:j<i\rangle=\mathbb{E}\langle\overline{x},b_{j}:j<i\rangle$. To conclude it suffices to invoke Corollary 29 with $\overline{y}$ an enumeration of the finite set $B\setminus\{b_{i}:i<\alpha\}$. ∎
We will use the following Lemma to remove the extra assumption from the above Corollary.
Lemma 34.
Let $(b_{i}:i<\alpha)$ be an ordinal-indexed sequence of elements of $\mathbb{E}_{*}$. Let $\overline{r}$ be a res-construction over $\mathbb{E}\langle b_{i}:i<\alpha\rangle$ and suppose that for all $\beta<\alpha$, $\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{r},b_{i}:i<\beta\rangle)\neq\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{r},b_{i}:i\leq\beta\rangle)$. Then for all $\beta<\alpha$, $\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i<\beta\rangle)\neq\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i\leq\beta\rangle)$. In particular $\mathbb{E}\langle b_{i}:i<\alpha\rangle$ is $\mathcal{O}_{*}$-res-constructible over $\mathbb{E}$.
Proof.
Note that if the hypotheses hold for $\alpha$, they also hold for any $\alpha^{\prime}<\alpha$. Therefore, by induction, we may assume the statement holds for all $\alpha^{\prime}<\alpha$. If $\alpha$ is a limit ordinal, then there is nothing to show. Suppose thus that $\alpha=\alpha^{\prime}+1$. Notice that $\overline{r}$ is in particular also an $\mathcal{O}_{*}$-res-construction over $\mathbb{E}\langle b_{i}:i<\alpha^{\prime}\rangle$, so by the inductive hypothesis $\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i<\beta\rangle)\neq\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i\leq\beta\rangle)$ for all $\beta<\alpha^{\prime}$ and we only need to show that $\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i<\alpha^{\prime}\rangle)\neq\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i\leq\alpha^{\prime}\rangle)$, i.e. that $b_{\alpha^{\prime}}$ is $\mathcal{O}_{*}$-residually generated over $\mathbb{E}\langle b_{i}:i<\alpha^{\prime}\rangle$. Notice that $b_{\alpha^{\prime}}$ is $\mathcal{O}_{*}$-residually generated over $\mathbb{E}\langle\overline{r},b_{i}:i<\alpha^{\prime}\rangle$ by hypothesis, so there is a finite $\overline{r}^{\prime}\subseteq\overline{r}$, a finite $F\subseteq\alpha^{\prime}$, and an $\mathbb{E}$-definable function $f$, such that $\operatorname{\mathbf{res}}(f(\overline{r}^{\prime},b_{F}))\notin\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{r},b_{i}:i<\alpha^{\prime}\rangle)$. Since then $\dim_{\operatorname{\mathrm{dcl}}}\big(\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{r}^{\prime},b_{i}:i<\alpha^{\prime}\rangle)/\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i<\alpha^{\prime}\rangle)\big)=|\overline{r}^{\prime}|$ and $\dim_{\operatorname{\mathrm{dcl}}}\big(\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{r}^{\prime},b_{i}:i\leq\alpha^{\prime}\rangle)/\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i<\alpha^{\prime}\rangle)\big)=|\overline{r}^{\prime}|+1$ it follows by a dimension argument that $\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i<\alpha^{\prime}\rangle)\neq\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i<\alpha\rangle)$. ∎
Theorem 35.
Suppose $\mathbb{E}_{*}\succ\mathbb{E}$ is $\mathcal{O}_{*}$-res-constructible over $\mathbb{E}$, $\operatorname{\mathbf{v}}_{\mathcal{O}_{*}}(\mathbb{E}_{*})$ contains no uncountable well-order and $\mathbb{E}_{*}\succ\mathbb{E}_{1}\succ\mathbb{E}$. Then $\mathbb{E}_{1}$ is $\mathcal{O}_{*}$-res-constructible.
Proof.
Let $S$ be a complement of $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}_{1})$ to a dcl-basis of $\mathbb{K}_{*}$. Set $\mathbb{E}_{2}:=\mathbb{E}_{1}\langle S\rangle$, $\mathbb{E}_{0}:=\mathbb{E}\langle S\rangle$ and let $B$ be a $\operatorname{\mathrm{dcl}}$-basis of $\mathbb{E}_{1}$ over $\mathbb{E}$. Notice that $\mathbb{E}_{*}$ is res-constructible over $\mathbb{E}_{0}$, $\mathbb{E}_{0}\prec\mathbb{E}_{2}\prec\mathbb{E}_{*}$, $\mathbb{E}_{2}=\mathbb{E}_{0}\langle B\rangle$, and $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}_{2})=\mathbb{K}_{*}$. By Corollary 33, we can find an ordinal-indexed enumeration $(b_{i}:i<\alpha)$ of $B$ such that $\operatorname{\mathbf{res}}(\mathbb{E}_{0}\langle b_{i}:i<\beta\rangle)\neq\operatorname{\mathbf{res}}(\mathbb{E}_{0}\langle b_{i}:i\leq\beta\rangle)$ for all $\beta<\alpha$. By Lemma 34, it follows that $\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i<\beta\rangle)\neq\operatorname{\mathbf{res}}(\mathbb{E}\langle b_{i}:i\leq\beta\rangle)$ for all $\beta<\alpha$. Thus one easily sees that $\mathbb{E}_{1}=\mathbb{E}\langle B\rangle$ is $\mathcal{O}_{*}$-res-constructible over $\mathbb{E}$. ∎
Theorem 36.
Suppose that $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{*},\mathcal{O}_{*})$ is res-constructible and such that:
  1. (1)
    $\operatorname{\mathbf{v}}(\mathbb{E}_{*})$ contains an uncountable well ordered subset;
  2. (2)
    $\mathbb{E}_{*}$ has uncountable $\operatorname{\mathrm{dcl}}_{T}$-dimension over $\mathbb{E}$.
Then there is $\mathbb{E}\prec\mathbb{E}_{1}\prec\mathbb{E}_{*}$ such that $\mathbb{E}_{1}$ is not $\mathcal{O}_{*}$-res-constructible and moreover $\operatorname{\mathbf{res}}(\mathbb{E}_{1})=\operatorname{\mathbf{res}}(\mathbb{E}_{*})$.
Proof.
By the hypothesis we can find $(s_{i}:i<\aleph_{1})$, $r$ and $(r_{i}:i<\aleph_{1})$ in $\mathbb{E}_{*}$ such that $\operatorname{\mathbf{res}}(r)$, $\operatorname{\mathbf{res}}(r_{i})_{i<\aleph_{1}}$, $\operatorname{\mathbf{res}}(s_{i})_{i<\kappa}$ are a $\operatorname{\mathrm{dcl}}_{T}$-basis of $\operatorname{\mathbf{res}}(\mathbb{E}_{*})$ over $\operatorname{\mathbf{res}}(\mathbb{E})$ and moreover $\mathbb{E}_{0}:=\mathbb{E}\langle s_{i}:i<\kappa\rangle$ contains a sequence $(x_{i}:i<\aleph_{1})$ such that $\operatorname{\mathbf{v}}(x_{i})>\operatorname{\mathbf{v}}(x_{j})$ for all $i<j$.
Set $\mathbb{F}_{i}^{\prime}:=\mathbb{E}_{0}\langle r+r_{j}x_{j}:j<i\rangle$. We claim that $\mathbb{F}:=\mathbb{F}^{\prime}_{\aleph_{1}}$ is not res-constructible. Suppose toward contradiction that it is: then there is a sequence $(\mathbb{F}_{i})_{i<\aleph_{1}}$, such that
  1. (1)
    $\mathbb{E}_{0}=\mathbb{F}_{0}\prec\mathbb{F}_{i}\prec\mathbb{F}_{j}\prec\mathbb{F}$ for all $0\leq i<j<\aleph_{1}$;
  2. (2)
    $\dim(\mathbb{F}_{i+1}/\mathbb{F}_{i})=1=\dim\big(\operatorname{\mathbf{res}}(\mathbb{F}_{i+1})/\operatorname{\mathbf{res}}(\mathbb{F}_{i})\big)$ and $\mathbb{F}_{i}=\bigcup_{j<i}\mathbb{F}_{j+1}$, for all $i<\aleph_{1}$.
Consider the two functions $f,g:\aleph_{1}\to\aleph_{1}$ defined respectively as
$f(\alpha):=\min\{\beta:\mathbb{F}_{\alpha}\subseteq\mathbb{F}^{\prime}_{\beta}\}\quad\text{and}\quad g(\alpha):=\min\{\beta:\mathbb{F}^{\prime}_{\alpha}\subseteq\mathbb{F}_{\beta}\}.$
Notice that these are both increasing and continuous.
It follows that there is some $0<\delta<\aleph_{1}$ which is a limit ordinal and a fixed point both of $f$ and $g$. For such $\delta$ we have $\mathbb{F}_{\delta}=\mathbb{F}^{\prime}_{\delta}$. Since by construction $\mathbb{F}$ is res-constructible over $\mathbb{F}_{\delta}$, by Lemma 20 all elements of $\mathbb{F}$ should be residually generated over $\mathbb{F}_{\delta}=\mathbb{F}^{\prime}_{\delta}$.
Notice that $r+x_{\delta}r_{\delta}\in r+x_{\beta}r_{\beta}+x_{\beta}\mathcal{O}_{*}$, thus since $r+x_{\delta}r_{\delta}$ cannot be weakly immediate over $\mathbb{F}_{\delta}$ (by Fact 12), it follows that there is $a\in\mathbb{F}_{\delta}$ such that $r+x_{\delta}r_{\delta}-a\in\bigcap_{\beta<\delta}x_{\beta}\mathfrak{o}_{*}$. Now there is a finite $F\subseteq\delta$, such that $a\in\mathbb{E}_{0}\langle r+r_{\beta}x_{\beta}:\beta\in F\rangle$. But now for all $\beta\in F$, $\operatorname{\mathbf{res}}(r_{\beta})=\operatorname{\mathbf{res}}((r+r_{\beta}x_{\beta}-a)/x_{\beta})\in\operatorname{\mathbf{res}}(\mathbb{E}_{0}\langle r+r_{\beta}x_{\beta}:\beta\in F\rangle)$ and on the other hand $\operatorname{\mathbf{res}}(r)=\operatorname{\mathbf{res}}(a)=\operatorname{\mathbf{res}}(r+r_{\beta}x_{\beta})\in\operatorname{\mathbf{res}}(\mathbb{E}_{0}\langle r+r_{\beta}x_{\beta}:\beta\in F\rangle)$ as well, thus since $\{\operatorname{\mathbf{res}}(r)\}\cup\{\operatorname{\mathbf{res}}(r_{\beta}):\beta\in F\}$ are independent over $\operatorname{\mathbf{res}}(\mathbb{E}_{0})$, it would follow that $\operatorname{\mathbf{res}}(\mathbb{E}_{0}\langle r+r_{\beta}x_{\beta}:\beta\in F\rangle)$ has dimension $|F|+1$ over $\operatorname{\mathbf{res}}(\mathbb{E}_{0})$ contradicting Fact 9. ∎
Proof of Theorem A.
The implication $(1)\Rightarrow(2)$ is Theorem 35, $(2)\Rightarrow(3)$ is trivial, and finally $(3)\Rightarrow(1)$ is Theorem 36. ∎
4. Application to Tressl’s pseudo-completions
We recall the definition of pseudo-completion of [9], and show how our results apply to [9, Problem 5.12].
Definition 37.
We call an elementary extension $\mathbb{K}\preceq\mathbb{K}^{\prime}\models T$, dense if $\mathbb{K}$ is dense in $\mathbb{K}^{\prime}$, we call it cofinal if $\mathbb{K}$ is cofinal in $\mathbb{K}^{\prime}$. Thus a Cauchy-completion of some $\mathbb{K}\models T$, can be defined as an elementary extension $\mathbb{K}_{*}\succeq\mathbb{K}$ which is maximal among the dense elementary extensions $\mathbb{K}^{\prime}\succeq\mathbb{K}$.
Remark 38.
In [9, Sec. 3], what we call Cauchy-completion is called just completion.
Lemma 39.
Suppose that $\mathbb{K}\preceq\mathbb{K}^{\prime}$ is a dense extension and suppose that $\mathbb{K}\preceq\mathbb{E}$ is a cofinal extension such that for all $x\in\mathbb{K}^{\prime}$, $\operatorname{\mathrm{tp}}(x/\mathbb{K})$ is realized in $\mathbb{E}$. Then there is a unique elementary embedding over $\mathbb{K}$ of $\mathbb{K}^{\prime}$ in $\mathbb{E}$.
Proof.
For the uniqueness, it suffices to show that for all $x\in\mathbb{K}^{\prime}\setminus\mathbb{K}$, $\operatorname{\mathrm{tp}}(x/\mathbb{K})$ has a unique realization in $\mathbb{E}$. If it was not the case and $x_{0},x_{1}\in\mathbb{E}$, were distinct realizations of $\operatorname{\mathrm{tp}}(x/\mathbb{K})$ then since $\mathbb{K}$ is dense in both $\mathbb{K}\langle x_{0}\rangle$ and $\mathbb{K}\langle x_{1}\rangle$ it would follow that $1/|x_{0}-x_{1}|>\mathbb{K}$ contradicting the cofinality of $\mathbb{K}$ in $\mathbb{E}$. To prove existence, let $(b_{i}:i<\alpha)$ be a $\operatorname{\mathrm{dcl}}_{T}$-basis of $\mathbb{K}^{\prime}$ over $\mathbb{K}$. To build an embedding of $\mathbb{K}$ in to $\mathbb{E}$ it suffices to show that for all $\beta<\alpha$, if $(c_{i}:i<\beta)$ is a tuple in $\mathbb{E}$ with $(b_{i}:i<\beta)\equiv_{\mathbb{K}}(c_{i}:i<\beta)$, then we can find $c_{\beta}$ in $\mathbb{E}$ such that $(b_{i}:i\leq\beta)\equiv_{\mathbb{K}}(c_{i}:i\leq\beta)$. This easily follows from the fact that $\operatorname{\mathrm{tp}}(b_{\beta}/\mathbb{K})\vdash\operatorname{\mathrm{tp}}(b_{\beta}/\mathbb{K}\langle b_{i}:i<\beta\rangle)$ because $\mathbb{K}^{\prime}$ is dense in $\mathbb{K}$. ∎
Corollary 40.
The following hold:
  1. (1)
    if $\mathbb{E}$ is a cofinal extension of $\mathbb{K}$, then there is a unique maximal dense extension of $\mathbb{K}$ in $\mathbb{E}$;
  2. (2)
    Cauchy-completions exist and are unique up to a unique isomorphism: if $\mathbb{K}\preceq\mathbb{K}^{\prime}$ and $\mathbb{K}\preceq\mathbb{K}^{\prime\prime}$ are Cauchy-completions of $\mathbb{K}$, then there is a unique isomorphism $\mathbb{K}^{\prime}\simeq\mathbb{K}^{\prime\prime}$ over $\mathbb{K}$.
Definition 41.
Let $\mathbb{E}\preceq\mathbb{U}\models T$ and $\mathcal{O}$ a $T$-convex subring of $\mathbb{U}$. We will call an extension $\mathbb{E}_{*}\preceq\mathbb{U}$ of $\mathbb{E}$ $\mathcal{O}$-residually dense (resp. $\mathcal{O}$-residually cofinal) if it is $\mathcal{O}$-res-constructible and moreover the extension $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})\preceq\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}_{*})$ is dense (resp. cofinal).
Lemma 42.
Suppose that $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ are $T$-convex valuation subrings of $\mathbb{E}_{1}$ and let $\mathbb{E}_{1}$ be $\mathcal{O}$-res-constructible over $\mathbb{E}$. If $\mathbb{E}_{1}$ is $\mathcal{O}$-residually cofinal over $\mathbb{E}$, or if $\mathcal{O}\subsetneq\mathcal{O}^{\prime}$, then $\mathbb{E}_{1}$ is $\mathcal{O}^{\prime}$-residually cofinal over $\mathbb{E}$.
Proof.
We first observe that if $\mathbb{K}\prec\mathbb{K}_{1}\models T$, $V$ is a non-trivial $T$-convex valuation ring of $\mathbb{K}_{1}$, and $\mathbb{K}_{1}$ is $V$-res-constructible over $\mathbb{K}$, then $\mathbb{K}$ is cofinal in $\mathbb{K}_{1}$. In fact if there was $t\in\mathbb{K}_{1}$ such that $t>\mathbb{K}$, then $t$ is $V$-purely valuational over $\mathbb{K}$ so by Fact 12 $t$ is not $V$-residually generated, contradicting Lemma 20.
Now let $\overline{r}$ be an $\mathcal{O}$-res-construction of $\mathbb{E}_{1}$ over $\mathbb{E}$, then by Lemma 23, it is also a $\mathcal{O}^{\prime}$-res construction and $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\overline{r})$ is a $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathcal{O})$-res-construction for $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})$ over $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$. In particular $\mathbb{E}_{1}$ is $\mathcal{O}^{\prime}$-res-constructible over $\mathbb{E}$ and $\mathbb{K}_{1}:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})$ is $V:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathcal{O})$-res-constructible over $\mathbb{K}:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$. Thus to conclude we only need to show that $\mathbb{K}$ is cofinal in $\mathbb{K}_{1}$.
If $\mathcal{O}^{\prime}\neq\mathcal{O}$, then $V\neq\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})$ and we are done by our first observation.
On the other hand if $\mathcal{O}^{\prime}=\mathcal{O}$ and $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$ is cofinal in $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}_{1})$, then the thesis is trivial. ∎
Remark 43.
In the situation of the Definition 41 above, if $\mathcal{O}$ and $\mathcal{O}^{\prime}$ are $T$-convex subrings of $\mathbb{U}$ such that $\mathcal{O}\cap\mathbb{E}=\mathcal{O}^{\prime}\cap\mathbb{E}$, then $\mathbb{E}_{*}\preceq\mathbb{U}$ is $\mathcal{O}$-residually dense (resp. $\mathcal{O}$-residually cofinal) over $\mathbb{E}$ if and only if it is $\mathcal{O}^{\prime}$-residually dense (resp. $\mathcal{O}^{\prime}$-residually cofinal). This is because if $\mathbb{E}_{*}$ is $\mathcal{O}$-residually cofinal over $\mathbb{E}$, then $\mathcal{O}\cap\mathbb{E}_{*}$ is the only $T$-convex valuation ring of $\mathbb{E}_{*}$ whose restriction to $\mathbb{E}$ is $\mathbb{E}\cap\mathcal{O}$.
Thus if $\mathcal{O}^{\prime\prime}$ is a $T$-convex valuation subring of $\mathbb{E}$, we will call an extension $\mathbb{E}_{*}$ of $\mathbb{E}$, $\mathcal{O}^{\prime\prime}$-residually dense (resp. $\mathcal{O}^{\prime\prime}$-residually cofinal) if it is $\mathcal{O}$-residually dense (resp. $\mathcal{O}$-residually cofinal) for some $T$-convex $\mathcal{O}$ in $\mathbb{U}$ with $\mathcal{O}\cap\mathbb{E}=\mathcal{O}^{\prime\prime}$.
Lemma 44.
If $\mathbb{E}\prec\mathbb{E}_{1}$ is a dense extension and $\mathcal{O}$ is a non-trivial $T$-convex valuation ring of $\mathbb{E}_{1}$, then $\mathbb{E}\not\subseteq\mathcal{O}$, $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})=\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}_{1})$, and $\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E})=\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E}_{1})$.
Proof.
The claim that $\mathbb{E}\not\subseteq\mathcal{O}$ is trivial, because if $\mathbb{E}\subseteq\mathcal{O}$, then by density of $\mathbb{E}$ in $\mathbb{E}_{1}$ it would follow that $\mathcal{O}=\mathbb{E}_{1}$.
To prove $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})=\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}_{1})$, suppose toward contradiction that there was an element $x$ of $\mathbb{E}_{1}\setminus\mathbb{E}$ such that $\operatorname{\mathbf{res}}_{\mathcal{O}}(x)\notin\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$. Then there would be no elements of $\mathbb{E}$ in $x+\mathfrak{o}$ whereas $(x+\mathfrak{o})\cap\mathbb{E}_{1}$ would be infinite, thus contradicting the density of $\mathbb{E}$ in $\mathbb{E}_{1}$.
Finally to see that $\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E})=\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E}_{1})$ notice that since $\mathbb{E}$ is dense in $\mathbb{E}_{1}$, then in particular for all $z\in\mathbb{E}_{1}^{>0}$, there is $c\in\mathbb{E}$ such that $c\geq z$ and |cz|/|z|<1|c-z|/|z|<1. ∎
Remark 45.
If $\overline{r}$ is a $\mathcal{O}$-res-construction over $\mathbb{E}$ and $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}_{1})=\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$, then $\overline{r}$ is a $\mathcal{O}$-res-construction over $\mathbb{E}_{1}$.
Lemma 46.
Let $\mathbb{E}\prec\mathbb{U}\models T$ and $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ be $T$-convex valuation subrings of $\mathbb{U}$ such that $\mathcal{O}\cap\mathbb{E}\neq\mathcal{O}^{\prime}\cap\mathbb{E}$. Suppose that $\mathbb{E}_{1}$ is a $\mathcal{O}^{\prime}$-residually dense extension of $\mathbb{E}$ with $\mathcal{O}^{\prime}$-res-construction $\overline{b}$ and $\mathbb{E}_{2}$ is a $\mathcal{O}$-res-constructible extension of $\mathbb{E}$ with $\mathcal{O}$-res-construction $\overline{c}$. Then the following hold:
  1. (1)
    $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})=\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}_{1})$ and $\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E}\cap(\mathcal{O}^{\prime}\setminus\mathfrak{o}^{\prime}))=\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E}_{1}\cap(\mathcal{O}^{\prime}\setminus\mathfrak{o}^{\prime}))$, in particular $\mathcal{O}\cap\mathbb{E}_{1}$ is the unique $T$-convex extension of $\mathcal{O}\cap\mathbb{E}$;
  2. (2)
    $\overline{c}$ is a $\mathcal{O}$-res-construction of $\operatorname{\mathrm{dcl}}(\mathbb{E}_{2}\cup\mathbb{E}_{1})$ over $\mathbb{E}_{1}$ and $\operatorname{\mathrm{dcl}}(\mathbb{E}_{2}\cup\mathbb{E}_{1})$ is $\mathcal{O}^{\prime}$-residually cofinal over $\mathbb{E}$ with $\mathcal{O}^{\prime}$-res-construction $(\overline{b},\overline{c})$;
  3. (3)
    $\operatorname{\mathrm{tp}}_{L}(\mathbb{E}_{1}/\mathbb{E})\vdash\operatorname{\mathrm{tp}}_{L}(\mathbb{E}_{1}/\mathbb{E}_{2})$.
Proof.
Throughout the proof let $V:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathcal{O})$ so that $\operatorname{\mathbf{res}}_{\mathcal{O}}=\operatorname{\mathbf{res}}_{V}\circ\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}$.
(1) Since $\mathbb{E}_{1}$ is $\mathcal{O}^{\prime}$-residually dense, $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})$ is a dense extension of $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$. Thus by Lemma 44 $\operatorname{\mathbf{res}}_{V}\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})=\operatorname{\mathbf{res}}_{V}\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$ and
$\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}_{1})=\operatorname{\mathbf{res}}_{V}\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})=\operatorname{\mathbf{res}}_{V}\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})=\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}).$
The statement about the value group is proved in the same way using that $\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E}_{1}\cap(\mathcal{O}^{\prime}\setminus\mathfrak{o}^{\prime}))=\operatorname{\mathbf{v}}_{V}(\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1}))$ and Lemma 44. The last clause is because in particular $\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E}_{1}\cap(\mathcal{O}^{\prime}\setminus\mathfrak{o}^{\prime}))=\operatorname{\mathbf{v}}_{V}(\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1}))$ and $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})=\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}_{1})$ imply that there is no $x\in\mathbb{E}_{1}$ such that $\mathcal{O}\cap\mathbb{E}<x<\mathbb{E}^{>\mathcal{O}}$.
(2) The first assertion is by point (1) and Remark 45. As for the second one, notice that by Lemma 23 it follows from the first assertion that $(\overline{b},\overline{c})$ is a $\mathcal{O}^{\prime}$-res-construction over $\mathbb{E}$, because $\overline{b}$ was a $\mathcal{O}^{\prime}$-res-construction over $\mathbb{E}$. Finally To see that $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$ is cofinal in $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\operatorname{\mathrm{dcl}}(\mathbb{E}_{2}\cup\mathbb{E}_{1}))$, recall that $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$ is dense in $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})$ by hypothesis and that $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})\prec\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\operatorname{\mathrm{dcl}}(\mathbb{E}_{2}\cup\mathbb{E}_{1}))$ is a $V$-res-constructible extension with $V\not\supseteq\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})$ and thus in particular it is cofinal by Lemma 42.
(3) Let $\overline{b}$ be a $\mathcal{O}^{\prime}$-res-construction of $\mathbb{E}_{1}$ over $\mathbb{E}$. Then it suffices to show that for any finite tuple $b^{\prime}_{0},\ldots b^{\prime}_{n}\subset\overline{b}$, $\operatorname{\mathrm{tp}}(b^{\prime}_{n}/\mathbb{E}\langle b^{\prime}_{0},\ldots b^{\prime}_{n-1}\rangle)\vdash\operatorname{\mathrm{tp}}(b^{\prime}_{n}/\mathbb{E}_{2}\langle b^{\prime}_{0},\ldots b^{\prime}_{n-1}\rangle)$.
For this notice that by Remark 45 and point (1), we have that $\mathbb{E}_{2}\langle b_{0}^{\prime},\ldots,b_{n-1}^{\prime}\rangle$ is $\mathcal{O}$-res-constructible over $\mathbb{E}\langle b_{0}^{\prime},\ldots b_{n-1}^{\prime}\rangle$. Thus by Lemma 20, every element of $\mathbb{E}_{2}\langle b_{0}^{\prime},\ldots,b_{n-1}^{\prime}\rangle$ is $\mathcal{O}$-residually generated over $\mathbb{E}\langle b_{0}^{\prime},\ldots,b_{n-1}^{\prime}\rangle$.
Now by point (1), $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}\langle b_{0}^{\prime},\ldots,b_{n}^{\prime}\rangle)=\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}\langle b_{0}^{\prime},\ldots,b_{n-1}^{\prime}\rangle)=\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$, thus $b_{n}^{\prime}$ is not $\mathcal{O}$-residually generated over $\mathbb{E}\langle b_{0}^{\prime},\ldots,b_{n-1}^{\prime}\rangle$. Moreover $\mathcal{O}\cap\mathbb{E}\langle b_{0}^{\prime},\ldots,b_{n}^{\prime}\rangle$ is the unique extension of $\mathcal{O}\cap\mathbb{E}\langle b_{0}^{\prime},\ldots,b_{n-1}^{\prime}\rangle$, so $\operatorname{\mathrm{tp}}_{L}(b_{n}^{\prime}/\mathbb{E}\langle b_{0}^{\prime},\ldots,b_{n-1}^{\prime}\rangle)$ implies the $L_{\mathrm{convex}}$-type of $b_{n}^{\prime}$ over $(\mathbb{E}\langle b_{0}^{\prime},\ldots,b_{n-1}^{\prime}\rangle,\mathcal{O}\cap\mathbb{E}\langle b_{0}^{\prime},\ldots,b_{n-1}^{\prime}\rangle)$. It follows that $\operatorname{\mathrm{tp}}_{L}(b_{n}^{\prime}/\mathbb{E}\langle b_{0}^{\prime},\ldots b_{n-1}^{\prime}\rangle)$ is not realized in $\mathbb{E}_{2}\langle b_{0}^{\prime},\ldots,b_{n-1}^{\prime}\rangle$ and we are done. ∎
Corollary 47.
Suppose that $\mathbb{E}\preceq\mathbb{U}\models T$ and let $\{\mathcal{O}_{i}:i\in I\}$ be a family of $T$-convex subrings of $\mathbb{U}$ such that for all $i\neq j$, $\mathcal{O}_{i}\cap\mathbb{E}\neq\mathcal{O}_{j}\cap\mathbb{E}$. Let also for each $i\in I$, $\mathbb{E}_{i}\preceq\mathbb{U}$ be a $\mathcal{O}_{i}$-residually dense extension of $\mathbb{E}$ with $\mathcal{O}_{i}$-res-construction $\overline{x}^{(i)}$. Then, setting $\mathcal{O}:=\bigcup_{i\in I}\mathcal{O}_{i}$:
  1. (1)
    $(x^{(i)}_{n}:i\in I,\,n<|\overline{x}^{(i)}|)$ is a $\mathcal{O}$-res-construction over $\mathbb{E}$ and $\operatorname{\mathrm{dcl}}(\bigcup_{i\in I}\mathbb{E}_{i})$ is $\mathcal{O}$-residually cofinal over $\mathbb{E}$;
  2. (2)
    $\bigcup_{i\in I}\operatorname{\mathrm{tp}}_{L}(\mathbb{E}_{i}/\mathbb{E})\vdash\operatorname{\mathrm{tp}}_{L}(\bigcup_{i\in I}\mathbb{E}_{i}/\mathbb{E})$;
Proof.
We first prove the statement with the additional hypothesis that $I$ is finite. If $I$ is empty there is nothing to show, otherwise let $j\in I$ be such that $\mathcal{O}_{j}\cap\mathbb{E}\supsetneq\mathcal{O}_{i}\cap\mathbb{E}$ for all $i\in I\setminus\{j\}$. By induction we may assume that the statement holds for $I^{\prime}=I\setminus\{j\}$. So $\bigcup_{i\in I\setminus\{j\}}\operatorname{\mathrm{tp}}(\mathbb{E}_{i}/\mathbb{E})\vdash\operatorname{\mathrm{tp}}\big((\bigcup_{i\in I\setminus\{j\}}\mathbb{E}_{i})/\mathbb{E}\big)$ and, with $\mathcal{O}^{\prime}:=\bigcup_{i\neq j}\mathcal{O}_{i}$, $\operatorname{\mathrm{dcl}}(\bigcup_{i\neq j}\mathbb{E}_{i})$ is $\mathcal{O}^{\prime}$-residually cofinal over $\mathbb{E}$ with $\mathcal{O}^{\prime}$-res-construction $(x^{(i)}_{n}:i\in I\setminus\{j\},\,n<|\overline{x}^{(i)}|)$.
But now by Lemma 46 (3) and (2) respectively, it follows that
$\operatorname{\mathrm{tp}}(\mathbb{E}_{j}/\mathbb{E})\cup\operatorname{\mathrm{tp}}\big((\bigcup_{i\in I\setminus\{j\}}\mathbb{E}_{i})/\mathbb{E}\big)\vdash\operatorname{\mathrm{tp}}(\bigcup_{i\in I}\mathbb{E}_{i}/\mathbb{E})$
and that $\operatorname{\mathrm{dcl}}(\bigcup_{i\in I}\mathbb{E}_{i})$ is $\mathcal{O}_{j}$-residually cofinal over $\mathbb{E}$ with $\mathcal{O}_{j}$-res-construction $(x^{(i)}_{n}:i\in I,\,n<|\overline{x}^{(i)}|)$.
Now suppose $I$ is infinite. Then (2) follows from a standard compactness argument. As for (1), observe that $(\operatorname{\mathbf{res}}_{\mathcal{O}}(x^{(i)}_{n}):i\in I,\,n<|\overline{x}^{(i)}|)$ is $\operatorname{\mathrm{dcl}}_{T}$-independent by Lemma 23 and by finite character of $\operatorname{\mathrm{dcl}}_{T}$. Similarly $\operatorname{\mathbf{res}}_{\mathcal{O}}(\operatorname{\mathrm{dcl}}(\bigcup_{i\in I}\mathbb{E}_{i}))$ must be a cofinal extension of $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$. ∎
Definition 48.
Let $\mathbb{E}\preceq\mathbb{U}\models T$ and $\mathcal{F}$ be a family of $T$-convex valuation subrings of $\mathbb{U}$ such that for all $\mathcal{O}\neq\mathcal{O}^{\prime}\in\mathcal{F}$, $\mathcal{O}\cap\mathbb{E}\neq\mathcal{O}^{\prime}\cap\mathbb{E}$. We will call $\mathcal{F}$-residually dense any extension $\mathbb{E}_{*}\preceq\mathbb{U}$ of $\mathbb{E}$ given as $\mathbb{E}_{*}=\operatorname{\mathrm{dcl}}\bigcup\{\mathbb{E}_{\mathcal{O}}:\mathcal{O}\in\mathcal{F}\}$ where $\mathbb{E}_{\mathcal{O}}$ is $\mathcal{O}$-residually dense over $\mathbb{E}$.
Remark 49.
In the situation of Definition 48, by Corollary 47, the isomorphism type of $\mathbb{E}_{*}$ over $\mathbb{E}$ only depends on the isomorphism types of the $\mathbb{E}_{\mathcal{O}}$ over $\mathbb{E}$, which in turn by Lemma 19 only depends on the isomorphism type of $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E}_{\mathcal{O}})$ over $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$.
Remark 50.
If $\mathbb{E}_{*}$ is a $\mathcal{F}$-residually dense extension of $\mathbb{E}$ and $\mathcal{O}$ is any $T$-convex valuation ring of $\mathbb{E}$, then $\operatorname{\mathrm{CH}}_{\mathbb{E}_{*}}(\mathcal{O})$ is the unique extension of $\mathcal{O}$ to a $T$-convex valuation ring of $\mathbb{E}_{*}$.
To see this, note first that the extension is unique if and only if $\mathbb{E}_{*}$ does not realise the type $\mathbb{E}<x<\mathbb{E}^{>\mathcal{O}}$, so we can reduce to the case where $\mathbb{E}\preceq\mathbb{E}_{*}$ is of finite $\operatorname{\mathrm{dcl}}_{T}$-dimension. Note next that if $\mathbb{E}\preceq\mathbb{E}_{\dagger}\preceq\mathbb{E}_{*}$, and $\mathcal{O}$ has a unique extension to $\mathbb{E}_{\dagger}$, say $\mathcal{O}_{\dagger}$, and $\mathcal{O}_{\dagger}$ has a unique extension to $\mathbb{E}_{*}$, then $\mathcal{O}$ must have a unique extension to $\mathbb{E}_{*}$. Thus, we can reduce further to the case where $\mathbb{E}\preceq\mathbb{E}_{*}$ is $1$-dimensional. In particular, we may assume $\mathbb{E}\preceq\mathbb{E}_{*}$ is $\mathcal{O}^{\prime}$-residually dense for some $\mathcal{O}^{\prime}\in\mathcal{F}$. If $\mathcal{O}^{\prime}\cap\mathbb{E}\subsetneq\mathcal{O}$, then Lemma 46 implies that the extension is $\mathcal{O}$-residually cofinal, whence the result follows. If they are equal, the result follows from Fact 12. Finally, if $\mathcal{O}^{\prime}\cap\mathbb{E}\supsetneq\mathcal{O}$, then any $\mathcal{O}^{\prime}$-residually dense extension cannot realize the cut $\mathbb{E}<x<\mathbb{E}^{>\mathcal{O}}$ because its realization would be $\mathcal{O}^{\prime\prime}$-residual where $\mathcal{O}^{\prime\prime}:=\{t\in\mathbb{U}:|t|<\mathbb{E}^{>\mathcal{O}}\}\subsetneq\mathcal{O}^{\prime}$.
Corollary 51.
Let $\mathbb{E}\preceq\mathbb{U}\models T$ and $\mathcal{F}$ be a family of $T$-convex valuation subrings of $\mathbb{U}$. Every $\mathcal{F}$-residually dense extension $\mathbb{E}\preceq\mathbb{E}_{*}\preceq\mathbb{U}$ has a $\operatorname{\mathrm{dcl}}_{T}$ basis $B$ over $\mathbb{E}$ such that for all $b\in B$, $\operatorname{\mathrm{tp}}(b/\mathbb{E})\vdash\operatorname{\mathrm{tp}}(b/\mathbb{E}\langle B\setminus\{b\}\rangle)$.
Definition 52 (Tressl, [9]).
Let $\mathbb{E}\models T$, $\mathcal{F}$ be a family of $T$-convex valuation rings of $\mathbb{E}$, $\mathbb{U}\succ\mathbb{E}$ be $|\mathbb{E}|^{+}$-saturated, and $\mathcal{F}^{\prime}:=\{\operatorname{\mathrm{CH}}_{\mathbb{U}}(\mathcal{O}):\mathcal{O}\in\mathcal{F}\}$. Let also for each $\mathcal{O}\in\mathcal{F}^{\prime}$, $\mathbb{E}_{\mathcal{O}}$ be a maximal $\mathcal{O}$-residually dense extension of $\mathbb{E}$ in $\mathbb{U}$. The $\mathcal{F}^{\prime}$-residually dense extension $\mathbb{E}_{*}:=\operatorname{\mathrm{dcl}}(\bigcup_{\mathcal{O}\in\mathcal{F}^{\prime}}\mathbb{E}_{\mathcal{O}})$ of $\mathbb{E}$ is called a pseudo-completion of $\mathbb{E}$ with respect to $\mathcal{F}$.
Corollary 53 (Tressl, 4.1 in [9]).
If $\mathcal{F}$ is a family of $T$-convex valuation rings of $\mathbb{E}$, then the pseudo-completion of $\mathbb{E}$ with respect to $\mathcal{F}$ is unique up to isomorphism over $\mathbb{E}$.
Proof.
Follows from Corollary 40 and Remark 49. ∎
Proposition 54.
Let $\mathbb{E}\prec\mathbb{U}\models T$ and $\mathcal{F}:=\{\mathcal{O}_{i}:0\leq i\leq n\}$ be strictly increasing sequence of $T$-convex valuation rings of $\mathbb{U}$ such that $\mathbb{E}\cap\mathcal{O}_{i}\neq\mathbb{E}\cap\mathcal{O}_{j}$ for all $i\neq j$ and $\mathcal{O}_{n}\supseteq\mathbb{E}$. For each $i\leq n$, let $\mathbb{E}_{i}$ be a $\mathcal{O}_{i}$-residually dense extension of $\mathbb{E}$ and let $\mathbb{E}_{*}:=\operatorname{\mathrm{dcl}}(\bigcup_{i\leq n}\mathbb{E}_{i})$ be $\mathcal{F}$-residually dense over $\mathbb{E}$. Moreover, for each cofinal extension $\mathbb{K}\preceq\mathbb{K}^{\prime}$ of models of $T$, denote by $d(\mathbb{K},\mathbb{K}^{\prime})$ the $\operatorname{\mathrm{dcl}}_{T}$-dimension of $\mathbb{K}^{\prime}$ over the maximal dense extension of $\mathbb{K}$ within $\mathbb{K}^{\prime}$. Then the following are equivalent:
  1. (1)
    if $\mathbb{E}\preceq\mathbb{E}_{\dagger}\preceq\mathbb{E}_{*}$ and for all $i\leq n$, $\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E}_{\dagger})$ contains $\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E}_{i})$, then $\mathbb{E}_{\dagger}$ is isomorphic to $\mathbb{E}_{*}$ over $\mathbb{E}$;
  2. (2)
    for all $0\leq i<n$, we have $d(\operatorname{\mathbf{res}}_{\mathcal{O}_{i+1}}(\mathbb{E}_{*}),\operatorname{\mathbf{res}}_{\mathcal{O}_{i+1}}(\mathbb{E}))\leq\aleph_{0}$ or $\operatorname{\mathbf{v}}(\mathbb{E}_{*},\mathcal{O}_{i})$ does not contain an uncountable well ordered subset.
Proof.
$(2)\Rightarrow(1)$. We induct on $n$. Notice that since $\mathbb{E}_{*}$ is a cofinal extension of $\mathbb{E}$, it follows that $\mathbb{E}_{*}\subset\mathcal{O}_{n}$. Then, the hypothesis $\operatorname{\mathbf{res}}_{\mathcal{O}_{n}}(\mathbb{E}_{n})\subset\operatorname{\mathbf{res}}_{\mathcal{O}_{n}}(\mathbb{E}_{\dagger})$ directly implies that $\mathbb{E}_{n}\subset\mathbb{E}_{\dagger}$. So if $n=0$ we are done, and we can assume $n>0$.
By repeatedly applying Lemma 46 (2), for all $0\leq i<n$ we know that $\operatorname{\mathrm{dcl}}(\mathbb{E}_{n},\mathbb{E}_{0},\mathbb{E}_{2},\cdots\mathbb{E}_{i})$ is $\mathcal{O}_{i}$-residually cofinal over $\mathbb{E}_{n}$ so in particular $\mathbb{E}_{*}$ is $\mathcal{O}_{n-1}$-res-constructible over $\mathbb{E}_{n}$. Now since $\dim_{\operatorname{\mathrm{dcl}}}(\mathbb{E}_{*}/\mathbb{E}_{n})\leq\aleph_{0}$ or $\operatorname{\mathbf{v}}_{\mathcal{O}_{n-1}}(\mathbb{E}_{*})$ does not contain uncountable well ordered sets, by Theorem 35, it follows that $\mathbb{E}_{\dagger}$ is also $\mathcal{O}_{n-1}$-res-constructible over $\mathbb{E}_{n}$.
Since by Lemma 19 the isomorphism type of $\mathbb{E}_{\dagger}$ over $\mathbb{E}_{n}$ only depends on the extension $\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E}_{\dagger})\succeq\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E}_{n})=\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E})$, it suffices to show that $\mathbb{E}_{\dagger}^{\prime}:=\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E}_{\dagger})$ is isomorphic to $\mathbb{E}_{*}^{\prime}:=\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E}_{*})$ over $\mathbb{E}^{\prime}:=\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E})$. Now let for all $i<n$, $\mathcal{O}_{n-1}^{\prime}:=\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathcal{O}_{i})$ and observe that for all $i<n$, $\operatorname{\mathbf{res}}_{\mathcal{O}_{i}^{\prime}}(\mathbb{E}_{*}^{\prime})=\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E}_{*})$, $\operatorname{\mathbf{res}}_{\mathcal{O}_{i}^{\prime}}(\mathbb{E}_{\dagger}^{\prime})=\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E}_{\dagger})$, and there is a natural inclusion $\operatorname{\mathbf{v}}_{\mathcal{O}_{i}^{\prime}}(\mathbb{E}_{*}^{\prime})\subseteq\operatorname{\mathbf{v}}_{\mathcal{O}_{i}}(\mathbb{E}_{*})$. Thus we can apply the inductive hypothesis to conclude that $\mathbb{E}_{*}^{\prime}$ is isomorphic to $\mathbb{E}_{\dagger}^{\prime}$ over $\mathbb{E}^{\prime}$.
$(\lnot 2)\Rightarrow(\lnot 1)$. Again we proceed by induction on $n$. Notice that if $n=0$, then (2) always holds, so we can assume $n>0$. Thus let $i<n$ be the maximum $i$ such that $d(\operatorname{\mathbf{res}}_{\mathcal{O}_{i+1}}(\mathbb{E}_{*}),\operatorname{\mathbf{res}}_{\mathcal{O}_{i+1}}(\mathbb{E}))>\aleph_{0}$ and $\operatorname{\mathbf{v}}(\mathbb{E}_{*},\mathcal{O}_{i})$ contains an uncountable well ordered subset. If i=n1i=n-1, then by Theorem 36, we can find $\mathbb{E}_{\dagger}$ as in (1) which is not $\mathcal{O}_{n-1}$-constructible over $\mathbb{E}_{n}$. But then, if $\phi:\mathbb{E}_{*}\rightarrow\mathbb{E}_{\dagger}$ is an isomorphism over $\mathbb{E}$, then $\phi$ must fix $\mathbb{E}_{n}$. So if $\overline{x}$ is a res-construction for $\mathbb{E}_{*}$ over $\mathbb{E}_{n}$, then $\phi(\overline{x})$ is a res-construction for $\mathbb{E}_{\dagger}$ over $\mathbb{E}_{n}$, a contradiction. If instead i<n1i<n-1, then $\dim_{\operatorname{\mathrm{dcl}}}(\mathbb{E}_{*}/\mathbb{E}_{n})\leq\aleph_{0}$ or $\operatorname{\mathbf{v}}(\mathbb{E}_{*},\mathcal{O}_{n-1})$ does not contain an uncountable well ordered set, so by Theorem 35 any $\mathbb{E}_{\dagger}$ as in (1) is $\mathcal{O}_{n-1}$-res-constructible over $\mathbb{E}_{n}$, and we can reduce to the inductive hypothesis. In fact we then have that the isomorphism type of $\mathbb{E}_{\dagger}$ over $\mathbb{E}_{n}$ is determined by the isomorphism type of $\mathbb{E}_{\dagger}^{\prime}:=\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E}_{\dagger})$ over $\mathbb{E}^{\prime}:=\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E}_{n})=\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E})$ and that, setting as before $\mathcal{O}_{i}^{\prime}:=\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathcal{O}_{i})$:
  • For any $0\leq i\leq n-1$, $\operatorname{\mathbf{res}}_{\mathcal{O}_{i}^{\prime}}(\mathbb{E}_{*}^{\prime})=\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E}_{*})$, $\operatorname{\mathbf{res}}_{\mathcal{O}_{i}^{\prime}}(\mathbb{E}^{\prime})=\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E})$ and the cokernel of the natural inclusion $\operatorname{\mathbf{v}}_{\mathcal{O}_{i}^{\prime}}(\mathbb{E}_{*}^{\prime})\subseteq\operatorname{\mathbf{v}}_{\mathcal{O}_{i}}(\mathbb{E}_{*})$ is $\operatorname{\mathbf{v}}_{\mathcal{O}_{n-1}}(\mathbb{E}_{*})$ which does not contain uncountable well orders by hypothesis, so $\operatorname{\mathbf{v}}_{\mathcal{O}_{i}^{\prime}}(\mathbb{E}_{*}^{\prime})$ contains an uncountable well order if and only if $\operatorname{\mathbf{v}}_{\mathcal{O}_{i}}(\mathbb{E}_{*})$ does.
Thus, by induction there exists some $\mathbb{E}_{\dagger}^{\prime}\prec\mathbb{E}_{*}^{\prime}$, such that for all $0\leq i\leq n-1$, $\operatorname{\mathbf{res}}_{\mathcal{O}_{i}^{\prime}}(\mathbb{E}_{i}^{\prime})\subset\operatorname{\mathbf{res}}_{\mathcal{O}_{i}^{\prime}}(\mathbb{E}_{\dagger}^{\prime})$, and which is non-isomorphic to $\mathbb{E}_{*}^{\prime}$ over $\mathbb{E}^{\prime}$. Let $\overline{x}$ be a tuple in $\mathbb{E}_{*}$ such that $\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\overline{x})$ is a $\operatorname{\mathrm{dcl}}$-basis for $\mathbb{E}_{\dagger}^{\prime}$ over $\mathbb{E}^{\prime}$, and let $\mathbb{E}_{\dagger}=\mathbb{E}_{n}\langle\overline{x}\rangle$. Then, for any $0\leq i\leq n-1$, $\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E}_{\dagger})=\operatorname{\mathbf{res}}_{\mathcal{O}_{i}^{\prime}}(\mathbb{E}_{\dagger}^{\prime})\supset\operatorname{\mathbf{res}}_{\mathcal{O}_{i}^{\prime}}(\mathbb{E}_{i}^{\prime})=\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E}_{i})$, and $\mathbb{E}_{n}\subset\operatorname{\mathbf{res}}_{\mathcal{O}_{n}}(\mathbb{E}_{\dagger})$, so $\mathbb{E}_{\dagger}$ satisfies the conditions of (1). Finally, $\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E}_{\dagger})=\mathbb{E}_{\dagger}^{\prime}$ is not isomorphic to $\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E}_{*})$ over $\operatorname{\mathbf{res}}_{\mathcal{O}_{n-1}}(\mathbb{E})$, so in particular $\mathbb{E}_{\dagger}$ is not isomorphic to $\mathbb{E}_{*}$ over $\mathbb{E}$. ∎
Remark 55.
Notice that in Proposition 54, the hypothesis that $\mathcal{O}_{n}\supseteq\mathbb{E}$ is not essential, as we can always assume that $\mathbb{E}_{n}$ is the trivial extension. Also notice that for all $i$, $\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E}_{i})$ is going to be a maximal dense extension of $\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E})$ in $\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E}_{*})=\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}\operatorname{\mathrm{dcl}}(\bigcup_{j\leq i}\mathbb{E}_{j})$. Thus
$d(\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E}_{*}),\operatorname{\mathbf{res}}_{\mathcal{O}_{i}}(\mathbb{E}))=\sum_{j<i}\dim_{\operatorname{\mathrm{dcl}}_{T}}(\mathbb{E}_{j}/\mathbb{E}).$
In particular in (2) of Proposition 54, if there is $i<n$ such that
$\dim_{\operatorname{\mathrm{dcl}}_{T}}(\operatorname{\mathbf{res}}_{\mathcal{O}_{i+1}}(\mathbb{E}_{*})/\operatorname{\mathbf{res}}_{\mathcal{O}_{i+1}}(\mathbb{E}))\leq\aleph_{0},$
then for all $j\leq i$, it follows that $d(\operatorname{\mathbf{res}}_{\mathcal{O}_{j+1}}(\mathbb{E}_{*}),\operatorname{\mathbf{res}}_{\mathcal{O}_{j+1}}(\mathbb{E}))\leq\aleph_{0}$.
It is natural to ask whether the conditions on $\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E}_{*})$ in point (2) of Proposition 54, can be substituted with conditions on $\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E})$, given that $\mathbb{E}_{*}$ is a very “narrow” extension of $\mathbb{E}$. We answer this conditionally to some properties of $T$.
Lemma 56.
Let $T$ be power bounded, $\mathbb{E}\prec\mathbb{U}\models T$, and $\mathcal{O},\mathcal{O}^{\prime}$ be $T$-convex valuation subrings of $\mathbb{U}$. If $\mathbb{E}_{*}\subseteq\mathbb{U}$ is a $\mathcal{O}^{\prime}$-residually dense extension of $\mathbb{E}$, then $\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E})\cong\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E}_{*})$.
Proof.
When $\mathcal{O}^{\prime}\subseteq\mathcal{O}$, then by Lemma 23, it follows that $\mathbb{E}_{*}$ is $\mathcal{O}$-res-constructible over $\mathbb{E}$ and the thesis follows from the rv-property (here Fact 13). Suppose instead $\mathcal{O}^{\prime}\supsetneq\mathcal{O}$. Let $\mathbb{K}:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$, $\mathbb{K}_{*}:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{*})$, $V:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathcal{O})$. The inclusion $\operatorname{\mathbf{v}}_{\mathcal{O}^{\prime}}(\mathbb{E})\subseteq\operatorname{\mathbf{v}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{*})$ induces a morphism of short exact sequences
[…]
Now since $\mathbb{K}_{*}$ is a dense extension of $\mathbb{K}:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$, by Lemma 44, the leftmost vertical map is an isomorphism. On the other hand by the case $\mathcal{O}^{\prime}\subseteq\mathcal{O}$, proved above, also the rightmost vertical map must be an isomorphism. So, by the short five lemma, the middle vertical map is an isomorphism. ∎
Corollary 57.
Let $T$ be power bounded, $\mathbb{E}\prec\mathbb{U}\models T$, and $\mathcal{F}$ be a family of $T$-convex valuation rings of $\mathbb{U}$ such that $\mathcal{O}\cap\mathbb{E}\neq\mathcal{O}^{\prime}\cap\mathbb{E}$ for all $\mathcal{O}\neq\mathcal{O}^{\prime}$ in $\mathcal{F}$. Suppose that $\mathbb{E}_{*}$ is a $\mathcal{F}$-residually dense extension of $\mathbb{E}$. Then $\operatorname{\mathbf{v}}(\mathbb{E},\mathcal{O})=\operatorname{\mathbf{v}}(\mathbb{E}_{*},\mathcal{O})$ for all $T$-convex valuation rings $\mathcal{O}$.
Proof.
By definition of $\mathcal{F}$-residually dense extension there are $\mathcal{O}$-residually dense extensions $\mathbb{E}_{\mathcal{O}}$ for all $\mathcal{O}\in\mathcal{F}$, such that $\mathbb{E}_{*}=\operatorname{\mathrm{dcl}}(\bigcup_{\mathcal{O}\in\mathcal{F}}\mathbb{E}_{\mathcal{O}})$. Notice that by the finite character of $\operatorname{\mathrm{dcl}}$ we can reduce to the case in which $\mathcal{F}=\{\mathcal{O}_{i}:i\leq n\}$ for some $n\in\omega$ and without loss of generality $\mathcal{O}_{i}\supsetneq\mathcal{O}_{j}$ for all $0\leq i<j\leq n$. Then set $\mathbb{E}_{i}:=\operatorname{\mathrm{dcl}}(\bigcup_{i<j}\mathbb{E}_{\mathcal{O}_{j}})$ for $0\leq i\leq n+1$ and notice that by Corollary 47, each $\mathbb{E}_{i+1}$ is a $\mathcal{O}_{i}$-residually dense extension of $\mathbb{E}_{i}$, moreover $\mathbb{E}_{0}=\mathbb{E}$ and $\mathbb{E}_{n+1}=\mathbb{E}_{*}$. At this point it suffices to apply Lemma 56 to conclude that for any $T$-convex valuation ring $\mathcal{O}^{\prime}$, $\operatorname{\mathbf{v}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{i})=\operatorname{\mathbf{v}}_{\mathcal{O}^{\prime}}(\mathbb{E})$. ∎
Remark 58.
When $T$ is not power-bounded, for a model $(\mathbb{E},\mathcal{O})\models T_{\mathrm{convex}}$, we have that $\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E})$ contains an uncountable well ordered subset if and only if $\mathbb{E}$ itself contains an uncountable well ordered subset.
In fact if $\mathbb{E}$ contains an uncountable well ordered set, then either $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$ or $\operatorname{\mathbf{v}}_{\mathcal{O}}(\mathbb{E})$ must contain an uncountable well ordered set (argue as in Lemma 22). On the other hand if $\mathbb{E}$ defines an exponential and $\operatorname{\mathbf{res}}_{\mathcal{O}}(\mathbb{E})$ contains an uncountable well ordered set $S$ and $d\in\mathbb{E}^{>\mathcal{O}}$, then $\operatorname{\mathbf{v}}_{\mathcal{O}}(\exp(Sd))$ is an uncountable well ordered set.
Thus, in view of Remark 58, when $T$ is not power-bounded, to see whether the condition that $\operatorname{\mathbf{v}}_{\mathcal{O}_{i}}(\mathbb{E})$ does not contain any uncountable well ordered set implies that $\operatorname{\mathbf{v}}_{\mathcal{O}_{i}}(\mathbb{E}_{*})$ does not contain any uncountable well ordered set we can just ask the same question about $\mathbb{E}$ and $\mathbb{E}_{*}$. This leads to consider the following property.
Definition 59.
We say that an o-minimal theory $T$ has property C if for every $\mathbb{E}\models T$, if $\mathbb{E}$ does not contain any uncountable well ordered set, then so does any 11-$\operatorname{\mathrm{dcl}}_{T}$-dimensional extension $\mathbb{E}\langle x\rangle$ of $\mathbb{E}$.
Lemma 60.
Let $\mathbb{E}\preceq\mathbb{U}\models T$, $\mathcal{F}$ be a family of $T$-convex subrings of $\mathbb{U}$ such that $\mathcal{O}\cap\mathbb{E}\neq\mathcal{O}^{\prime}\cap\mathbb{E}$ for all $\mathcal{O}\neq\mathcal{O}^{\prime}$ in $\mathcal{F}$, and let $\mathbb{E}_{*}\preceq\mathbb{U}$ be $\mathcal{F}$-residually dense over $\mathbb{E}$. Suppose furthermore that $T$ has property C. If $\mathbb{E}_{*}$ contains an uncountable well-ordered set, then so does $\mathbb{E}$.
Proof.
Suppose toward contradiction that $\mathbb{E}_{*}$ contains an uncountable well-ordered set and $\mathbb{E}$ does not.
By Corollary 51 we can find a $\operatorname{\mathrm{dcl}}_{T}$-basis $B$ of $\mathbb{E}_{*}$ over $\mathbb{E}$ such that for all $b\in B$, $\operatorname{\mathrm{tp}}(b/\mathbb{E})\vdash\operatorname{\mathrm{tp}}(b/\mathbb{E}\langle B\setminus\{b\}\rangle)$.
Let $(y_{i})_{i<\aleph_{1}}$ be a sequence in $\mathbb{E}_{*}$ such that $y_{i}$ is strictly increasing. Since $\mathbb{E}$ does not contain an uncountable well ordered set, there must be $i_{0}$ such that for all $i>i_{0}$, $\operatorname{\mathrm{tp}}(y_{i}/\mathbb{E})=\operatorname{\mathrm{tp}}(y_{i_{0}}/\mathbb{E})$.
For each $i$, write $y_{i}:=f_{i}(\overline{b}^{(i)})$ for some tuple $\overline{b}^{(i)}$ from $B$ and some $\mathbb{E}$-definable function $f_{i}$. The orthogonality property of $B$ entails that that there is a non-empty $B_{0}\subseteq B$ such that for every $\mathbb{E}$-definable $f$, if $\operatorname{\mathrm{tp}}(f(\overline{b})/\mathbb{E})=\operatorname{\mathrm{tp}}(y_{i_{0}}/\mathbb{E})$ for some tuple $\overline{b}$ from $B$, then $\overline{b}\supseteq B_{0}$. Thus all tuples $\overline{b}^{(i)}$ for $i>i_{0}$ must contain $B_{0}$.
Since $B_{0}$ is finite and thus by property C, $\mathbb{E}\langle B_{0}\rangle$ still does not contain an uncountable well-ordered set, we can repeat the argument on the sequence $(y_{i})_{i_{0}\leq i<\aleph_{1}}$, but this time over $\mathbb{E}\langle B_{0}\rangle$, instead of $\mathbb{E}$.
Thus by induction we get an increasing sequence $(i_{n}:n<\omega)$ in $\aleph_{1}$, such that for $i\geq i_{n}$, $\overline{b}^{(i)}$ must contain at least $n$ distinct elements from $B$. Since $\{i_{n}:n<\omega\}$ is bounded in $\aleph_{1}$ we get a contradiction. ∎
We will discuss further which theories have property C in a later version of this paper.